### v3

# Convergence Rate for the Last Iterate of Stochastic Gradient Descent Schemes

Abstract.
We study the convergence rate for the last iterate of stochastic gradient descent (SGD) and stochastic heavy ball (SHB) in the parametric setting when the objective function $F$ is globally convex or non-convex whose gradient is $\gamma$-Hölder. Using only discrete Gronwall’s inequality without Robbins-Siegmund theorem, we recover results for both SGD and SHB: $\min_{s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ for non-convex objectives and $F(w_{\tau\wedge t})-F_{*}=o(t^{2\gamma/(1+\gamma)\cdot\max(p-1,-2p+1)-\epsilon})$ for $\beta\in(0,1)$, $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$, and $\min_{s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ for convex objectives $F$ whose minimum is $F_{*}$. In addition, we proved that SHB with constant momentum parameter $\beta\in(0,1)$ attains a convergence rate of $F(w_{t})-F_{*}=O(t^{\max(p-1,-2p+1)}\log^{2}\frac{t}{\delta})$ with probability at least $1-\delta$ when $F$ is convex and $\gamma=1$ and step size $\alpha_{t}=\Theta(t^{-p})$ with $p\in(\frac{1}{2},1)$.
Key words and phrases:
stochastic optimization, stochastic gradient descent, stochastic heavy ball, almost sure convergence, convergence rate with high probability
1. Introduction
We study the almost sure convergence rate for the last iterate of Stochastic Gradient Descent (SGD) and Stochastic Heavy Ball (SHB) schemes associated to solving an unconstrained optimization problem involving a cost function defined on a subset of $\mathbb{R}^{d}$. We consider the application of these algorithms in the following context as in [2]:
$w_{t+1}-w_{t}=-\alpha_{t}\,\nabla\ell(Z_{t},w_{t})+\beta(w_{t}-w_{t-1})~,~\beta\in[0,1)$ (1.1)
where $w_{t}$ is in $\mathbb{R}^{d}$, $Z_{t}\in\mathbb{R}^{N}$ is an i.i.d process with finite variance sampled from a continuous distribution with density $\rho$ independent of $\mathscr{F}_{t}=\sigma(w_{s}:s\leq t)$, $\alpha_{t}$ is the (non-random) step size, and $\ell$ is an estimator of a deterministic cost function $F:\pazocal{W}\rightarrow[0,\infty)$ defined by $F(w)=\mathbb{E}_{\rho}[\ell(Z,w)]$. When $\beta=0$, (1.1) is the SGD and can be rewritten as
$w_{t+1}-w_{t}=-\alpha_{t}\nabla F(w_{t})+\alpha_{t}\,\delta m_{t}~~~,~~~\delta m_{t}=\nabla F(w_{t})-\nabla\ell(Z_{t},w_{t}).$ (1.2)
Algorithms of the form (1.2) have been studied extensively as part of a field called stochastic approximation ([10], [13]). However, the almost sure convergence rate for the last iterate of SGD is studied only recently. The most recent work in the parametric setting are done by Weissmann et al in [15]. Weissmann et al. considers SGD and SHB where properties of the objective function $F$ is only known locally, i.e. properties apply in some neighborhood of a point [15, assumption 2.1.(ii)]. Under the local gradient domination [15, definition 2.2] $\|\nabla F(w_{t})\|^{2}\geq F(w_{t})-F_{*}$, Weissmann et al proved the existence of neighborhoods $\pazocal{U}_{1},\pazocal{U}$ of critical points in the parameter space $\pazocal{W}$ for which if $w_{1}\in\pazocal{U}_{1}$, then the event $\Omega{U}=\{\pazocal{U}\,\text{is absorbing for}\,w_{t}\}$ is highly probable and $(F(w_{t})-F_{*})\mathbbm{1}_{\Omega{U}}=o(t^{-1+\epsilon})$ [15, table 1, theorem 5.1]. The set $\pazocal{U}$ is defined as a neighborhood of the set of minimizers with a fixed radius. We remark that Weissmann et al also examines the case when global properties of $F$ are known [15, table 1].
When global properties of the objective function are known, there are other recent works as well. Liu and Yuan proved almost sure convergence rates for SGD, SHB, and SNAG (Stochastic Nesterov’s Accelerated Gradient) by assuming that the objective function is either globally non-convex, convex, or strongly convex [9]. Liu and Yuan produced the rate $F(w_{t})-F_{*}=o(t^{-1+\epsilon})$ for strongly convex $F$ and $\min_{1\leq i<t-1}\|\nabla F(w_{i})\|^{2}=o(\sum_{i=1}^{t-1}\alpha_{i})^{-1}$ for non-convex $F$ [9, theorem 6, theorem 8]. The strongly convex rate is optimal according to Agarwal et al [1]. For a convex objective function $F$, Liu and Yuan proved a rate of $F(w_{t})-F_{*}=o(t^{-\frac{1}{3}+\epsilon})$, which is not as close to the optimal rate of $t^{-1/2}$ in expectation [9, remark 14]. Liu and Yuan used the Robbins-Siegmund theorem ([13, theorem 1], [9, proposition 2]) to prove their convergence rate results. We remark that both SGD and SHB produce the same rate according to Liu and Yuan, i.e. the momentum parameter $\beta$ does not affect the convergence rate [9, theorem 8, theorem 13].
Intuitively, the momentum parameter $\beta$ in (1.1) serves to counter slowdown when the gradient is near zero. As such, it makes sense to adjust $\beta$ depending on the gradient and step sizes. Sebbouh et al studied such a case in [14]. In particular, they proved that the last iterate of SHB-IMA (Iterate Moving Average) can achieve $o(t^{-1+\epsilon})$ rate in the convex setting when $\beta$ is allowed to vary as a function of $t$ in the ‘overparametrized’ case [14, table 1, corollary 14, definition 3]. On the other hand, in the usual parametrized case, given step size of order $O(t^{-\frac{1}{2}-\epsilon})$, Sebbouh et al achieved $o(t^{-\frac{1}{2}+\epsilon})$ rate [14, corollary 17]. Similar to Liu and Yuan’s work [9], Sebbouh et al used the Robbins-Siegmund theorem [14, lemma 6] to prove these rates.
In another work [8], Lei, Shi, and Guo obtained the last iterate almost sure convergence rate of $o(t^{\max(p-1,1-p(\gamma+1))+\epsilon})$ given step size $O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ for a convex objective function, whose gradient is $\gamma$-Hölder with SGD in the non-parametric setting [8, theorem 6]. Moreover, in [8, corollary 11], Lei, Shi, and Guo proved that SGD converges at rate $F(w_{t+1})-F_{*}=O(t^{\max(p-1,1-p(\gamma+1))}\,\log^{2}(t/\delta))$ with probability at least $1-\delta$ if $p\in(\frac{1}{1+\gamma},1)$. In particular, when $p=2/3$ and $\gamma=1$, their work matches Liu and Yuan’s with $O(t^{-\frac{1}{3}+\epsilon})$ [8, theorem 6, corollary 11]. Interestingly, Lei, Shi, and Guo did not use Robbins-Siegmund theorem. However, their approach is similar to that of Robbins and Siegmund, where a supermartingale is constructed and the convergence of the iterates is implied by that of the supermartingale [8, section 4.2].
Another recent work regarding almost sure convergence rate for SGD without employing the Robbins-Siegmund theorem, is done by Nguyen et al [11]. The authors used a weaker, super-martingale convergence theorem [11, lemma 5] to prove $w_{t}\rightarrow w_{*}$ a.s.-$\mathbb{P}$ and $\mathbb{E}\|w_{t}-w_{*}\|^{2}=o(t^{-1})$ in the strong convex case [11, thorem 1,2]. Such a rate is indeed optimal according to [1] since if $F$ is convex with $\nabla F$ Lipschitz, by lemma C.4, we get $F(w_{t})-F_{*}\lesssim\|w_{t}-w_{*}\|^{2}=o(t^{-1})$.
1.1. Contribution
As remarked in Liu and Yuan’s paper [9, remark 14], the rate $O(t^{-\frac{1}{3}+\epsilon})$ for the convex case is not close to the optimal rate in expectation $O(t^{-\frac{1}{2}})$ proved by Agarwal et al in [1]. Motivated by this remark, we study the almost sure convergence for the last iterate of SGD and SHB where the objective function $F$ is globally convex and non-convex with $\nabla F$ being $(\gamma,L)$-Hölder. With these conditions, our contribution is as follows:
  1. 1.
    Instead of relying on Robbins-Siegmund, nor any stochastic approximation theory, we use only Gronwall’s inequality without using any martingale convergence theorem to prove that $\min_{0\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ and $F(w_{t})-F_{*}=o(t^{(2\gamma/(1+\gamma))\max(p-1,1-(1+\gamma)p)-\epsilon})$, as well as $\min_{s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ for non-convex and convex objectives respectively given step size of $\alpha_{t}=O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ using similar assumptions as in [9], except that $\nabla F$ is assumed to be $\gamma$-Hölder instead of Lipschitz. By using non-martingale convergence theory, our argument is less technical. We also present insights from our study in section 2.2 and 2.3.
  2. 2.
    We proved that with probability at least $1-\delta$, SHB (1.1) with a fixed momentum parameter $\beta\in(0,1)$ converges at rate $F(w_{t})-F_{*}=O(t^{\max(p-1,-2p+1)}\log^{2}\frac{t}{\delta})$ when the gradient of the objective function $\nabla F$ is Lipschitz and $F$ is convex. Our result is similar to [8] and is consistent with our own findings (see table 1). While our result is restricted to the case when $\beta$ is fixed, such a result provides baseline SHB performance which can be compared to other algorithms, e.g. SHB-IMA (Iterated Moving Average). There may be arguments to use baseline (simpler) SHB variant to save computation time. To the author’s knowledge, this is the first high probability convergence rate result for the SHB.
  3. 3.
    We use a slightly weaker assumption for $\nabla F$ compared to Liu and Yuan [9, assumption 4] that it is only $\gamma$-Hölder with $\gamma\in(0,1]$, following Lei, Shi, and Guo [8, assumption 1]. In this manner, we provide a unified analysis for ‘smooth’ objective functions.
1.2. Acknowledgment
I would like to thank Prof. Sunder Sethuraman, for discussions and support throughout the writing process of this manuscript.
Algo Cost Function w/ $(\gamma,L)$-smoothness Step Size Decay Rate $p$ in $O(t^{-p})$ Convergence Rate for $F(w_{t})-F_{*}$ (convex case) or $\|\nabla F(w_{t})\|^{2}$ (non-convex) Statement
SGD Convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $o(t^{\max(p-1,1-(\gamma+1)p)-\epsilon})$ theorem 2.7, [8, theorem 6].
$(\frac{1}{2},1)$ $O(t^{\max(p-1,1-(\gamma+1)p)}\log^{2}\frac{t}{\delta})$ with prob. $1-\delta$ [8, corollary 11].
Convex, $\gamma=1$ $(\frac{2}{3},1)$ $o(t^{p-1})$ [9, theorem 13].
Non-convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $\min_{0\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ corollary 3.3.
Non-convex, $\gamma=1$ $(\frac{1}{2},1)$ $\min_{1\leq i\leq t-1}\|\nabla F(w_{t})\|^{2}=o\left(\frac{1}{\sum_{i=1}^{t-1}\alpha_{i}}\right)$ [9, theorem 6].
SHB Convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $o(t^{\frac{2\gamma}{1+\gamma}\max(p-1,1-(\gamma+1)p)-\epsilon})$ theorem 2.7.
Convex, $\gamma=1$ $(\frac{2}{3},1)$ $o(t^{p-1})$ [9, theorem 13].
$(\frac{1}{2},1)$ $O(t^{\max(p-1,1-2p)}\log^{2}\frac{t}{\delta})$ w/ prob. $1-\delta$ theorem 2.10.
Non-convex, $\gamma\in(0,1]$ $(\frac{1}{1+\gamma},1)$ $\min_{1\leq s\leq t}\|\nabla F(w_{t})\|^{2}=o(t^{p-1})$ corollary 3.6.
Non-convex, $\gamma=1$ $(\frac{1}{2},1)$ $\min_{1\leq i\leq t-1}\|\nabla F(w_{t})\|^{2}=o\left(\frac{1}{\sum_{i=1}^{t-1}\alpha_{i}}\right)$ [9, theorem 8].
SHB-IMA Convex, $\gamma=1$ $\frac{1}{2}-\epsilon$ $o(t^{-\frac{1}{2}+\epsilon})$ [14, corollary 17].
Table 1. Summary of Results.
2. Mathematical Model and Main Results
2.1. Notation and Assumptions
We list notations, assumptions, results, and main mathematical arguments that we use in our work for convenience of the reader.
Symbol Description
$\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ The set of non-negative integers.
$\mathbb{R}_{+}$ The open interval $(0,\infty)$, i.e. the set of strictly positive real numbers.
$\lesssim$ $a\lesssim b$ means that there exists $c>0$ such that $a\leq c\cdot b$.
$\|\cdot\|$ The standard Euclidean norm in $\mathbb{R}^{d}$.
$f(t)=o(g(t))$ $f(t)/g(t)\rightarrow 0$ as $t\rightarrow\infty$.
$f(t)=O(g(t))$ $\exists\,c>0$ and $T>0$ such that $f(t)\leq c\,g(t)$ for all $t>T$.
$f(t)=\Theta(g(t))$ $f(t)=O(g(t))$ and $g(t)=O(f(t))$.
$(\Omega,\mathscr{F},\mathbb{P})$ The probability space for the paired process $(Z_{t},w_{t})$ in (1.1).
$\mathscr{F}_{t}$ The filtration $\{\mathscr{F}_{t}:t\in\mathbb{N}\}$ where $\mathscr{F}_{t}=\sigma\{w_{s}:s\leq t\}$.
$\mathscr{F}_{\infty}$ The $\sigma$-algebra $\bigcup_{t\in\mathbb{N}}\mathscr{F}_{t}$.
$m\mathscr{F}_{t}$ The set of measurable functions w.r.t $\mathscr{F}_{t}$ for $t\in\mathbb{N}\cup\{\infty\}$.
$E_{n}$ eventually The event $\bigcup_{m=1}^{\infty}\bigcap_{n\geq m}E_{n}$ ($\exists m>0$ such that $E_{n}$ occurs for all $n\geq m$).
$\ell(z,w)$ The estimator $\ell:\pazocal{Z}\times\pazocal{W}\rightarrow[0,\infty)$ of the objective function.
$Z_{t}$ An i.i.d sequence of random vectors in $\pazocal{Z}\subset\mathbb{R}^{N}$ that makes up the stochastic gradient estimator $\nabla\ell(Z_{t},w_{t})$.
$\rho$ The density function for $Z_{t}$.
$F$ The objective function $F:\pazocal{W}\rightarrow[0,\infty)$ defined by $F(w)=\mathbb{E}_{\rho}[\ell(Z_{t},w)]$.
$\delta m_{t}$ The injected noise in stochastic algorithms, defined in (1.2).
Table 2. Summary of Notations.
Assumption 2.1.
Let $\pazocal{Z}\subset\mathbb{R}^{n}$, $\pazocal{W}\subset\mathbb{R}^{d}$ and $\rho$ be a probability density on $\pazocal{Z}$. The function $\ell:\pazocal{Z}\times\pazocal{W}\rightarrow\mathbb{R}_{+}$ satisfies:
  1. 1.
    $\ell(z,\cdot)$ is $(\gamma,L)$-smooth, i.e. $\ell(z,\cdot)$ is differentiable and $\nabla\ell(z,\cdot)$ is $(\gamma,L)$-Hölder:
    $\exists\gamma\in(0,1]~\text{such that}~\|\nabla\ell(z,u)-\nabla\ell(z,v)\|\leq L\,\|u-v\|^{\gamma}~~~\forall\,u,v\in\pazocal{W},\,z\in\pazocal{Z}.$
  2. 2.
    $\nabla\ell(\cdot,w)\in L^{1}(\pazocal{Z},\mathscr{B},\rho\,dz)$ for all $w\in\pazocal{W}$ where $\mathscr{B}$ is the Borel $\sigma$-algebra in $\mathbb{R}^{N}$.
  3. 3.
    $F(w)=\mathbb{E}_{\rho}[\ell(Z,w)]$ is bounded below by $F_{*}=\inf_{w\in\pazocal{W}}F(w)>$-$\infty$ and a global minimizer $w_{*}\in\pazocal{W}$ exists, i.e. $F(w_{*})=F_{*}$.
The assumption that $\nabla\ell$ is only $\gamma$-Hölder is used in [8, assumption 1], whose application is also mentioned in [8]. This is a weakening of the $\nabla\ell$ being Lipschitz used in [9, assumption 1]. However, thanks to a slight modification in the proof of Garrigos and Gower [4, lemma 2.25] (see proposition C.2), any function $f$ that is $(\gamma,L)$-smooth still enjoys the property
$f(y)\leq f(x)+\langle\nabla f(x),\,y-x\rangle+\frac{L}{1+\gamma}\|y-x\|^{1+\gamma}~~~\,\forall x,y\in\mathbb{R}^{d}.$ (2.1)
Now, observe that $\mathbb{E}_{\rho}[\nabla\ell(Z,w)]$ exists due to the second property in assumption 2.1. Such a property can be satisfied for example when $\pazocal{Z}$ is bounded or if $Z$ is Gaussian. Moreover, assuming such property is natural; in the gradient descent (GD) scheme
$w_{t+1}-w_{t}=-\alpha_{t}\,\frac{1}{n}\sum_{k=1}^{n}\nabla\ell(Z_{k},w_{t})$ (2.2)
such an assumption implies convergence almost surely of the empirical average on the right hand side to $\mathbb{E}_{\rho}[\nabla\ell(Z,w_{t})]=\nabla\mathbb{E}_{\rho}[\ell(Z,w_{t})]=\nabla F(w_{t})$ as the number of data points $n\uparrow\infty$. This suggests that (2.2) approximates the deterministic GD [4, algorithm 3.2]. In addition, there are important consequences. The first consequence is due to proposition C.1: $F$ inherits the properties in assumption 2.1 from the function $\ell(z,\cdot)$. Secondly, if the joint distribution of $(Z_{t},w_{t})$ is the product measure $\rho\,dz\otimes\mu_{t}$, then
$\int\mathbbm{1}_{\{w_{t}\in A\}}\,\mathbb{E}[\nabla\ell(Z_{t},w_{t})\,|\,\mathscr{F}_{t}]\,d\mathbb{P}$ $=\int_{A}\int_{\pazocal{Z}}\nabla\ell(z,w)\,\rho(z)\,dz\,\mu_{t}(dw)$ (2.3)
$=\int_{A}\mathbb{E}_{\rho}[\nabla\ell(Z,w)]\,\mu_{t}(dw)=\int_{A}\nabla\mathbb{E}_{\rho}[\ell(Z,w)]\,\mu_{t}(dw)$
$=\int\mathbbm{1}_{\{w_{t}\in A\}}\,\nabla F(w_{t})\,d\mathbb{P}.$
Equation (2.3) implies $\mathbb{E}[\nabla\ell(Z_{t},w_{t})\,|\,\mathscr{F}_{t}]=\nabla F(w_{t})$ and that the estimator $\nabla\ell(Z,w_{t})$ is unbiased (w.r.t $\rho$). That $\nabla F(w_{t})$ can be accessed through the estimator $\nabla\ell(Z,w_{t})$ is also assumed in [9, the sentence right above assumption 4].
Assumption 2.2.
Using the same notation as in assumption 2.1, there exists $A,B,C\in\mathbb{R}_{+}$ such that $\mathbb{E}[\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}\,|\,\mathscr{F}_{t}]\leq A(F(w_{t})-F_{*})+B\,\|\nabla F(w_{t})\|^{1+\gamma}+C$ for all $t>0$ a.s.-$\mathbb{P}$.
The above assumption is called the ABC condition, originally proposed by Khaled and Richtárik in [6], and is used by Liu and Yuan in [9, assumption 4] when $F$ is $(1,L)$-smooth, i.e. $\gamma=1$. It is said to be “the weakest assumption” for analysis of SGD in the non-convex setting [9, remark 1]. By lemma C.5, the first bounding term is a consequence of $(\gamma,L)$-smoothness and convexity:
$\int\mathbbm{1}_{\{w_{t}\in A\}}$ $\mathbb{E}[\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}\,|\,\mathscr{F}_{t}]\,d\mathbb{P}=\int_{A}\mathbb{E}_{\rho}\|\nabla\ell(z,w)\|^{1+\gamma}\,\mu_{t}(dw)$
$\leq\int_{A}\bigg\{c_{1}(\beta,\gamma)(F(w)-F_{*})+c_{2}(\beta,\gamma)+c_{3}(\beta,\gamma)\,\mathbb{E}_{\rho}[\|\nabla\ell(Z,w_{*})\|^{1+\gamma}]\bigg\}\,\mu_{t}(dw).$
Note that the last two terms in the integrand are constants and that the existence of a global minimizer $w_{*}\in\pazocal{W}$ is required (satisfied by assumption 2.1). When $\gamma=1$ and the gradient estimator is unbiased, i.e. $\mathbb{E}_{\rho}[\nabla\ell(Z,w)]=F(w)$, the variance is
$\mathbb{E}_{\rho}\|\nabla\ell(Z,w)-\nabla F(w)\|^{2}$ $\leq\mathbb{E}_{\rho}\|\nabla\ell(Z,w)\|^{2}-2\,\mathbb{E}_{\rho}[\langle\nabla\ell(Z,w),\,\nabla F(w)\rangle]+\|\nabla F(w)\|^{2}$
$=\mathbb{E}_{\rho}\|\nabla\ell(Z,w)\|^{2}-\|\nabla F(w)\|^{2}.$
Therefore, if the variance is bounded by $\sigma^{2}<\infty$, then $\mathbb{E}_{\rho}\|\nabla\ell(Z,w)\|^{2}\leq\|\nabla F(w)\|^{2}+\sigma^{2}$. This justifies the second term in the ABC assumption. For $\gamma<1$, assuming that the gradient estimator has variance bounded by $\sigma^{2}$, we end up with
$\mathbb{E}_{\rho}\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}\leq 2^{1+\gamma}(\|\nabla F(w_{t})\|^{1+\gamma}+\mathbb{E}_{\rho}\|\delta m_{t}\|^{1+\gamma})\leq B\,\|\nabla F(w_{t})\|^{1+\gamma}+C\,\sigma^{1+\gamma}.$
In fact, with a simple estimate $a^{1+\gamma}\leq(a+1)^{1+\gamma}\leq(a+1)^{2}\leq 2a^{2}+2$, setting $a=\|\nabla F(w_{t})\|$ gives
$\mathbb{E}[\|\nabla\ell(Z_{t},w_{t})\|^{1+\gamma}\,|\,\mathscr{F}_{t}]\leq 2B\,\|\nabla F(w_{t})\|^{2}+(C\,\sigma^{1+\gamma}+2)~~~\text{a.s.-}\mathbb{P}.$ (2.4)
which has the form of the ABC condition in [9, assumption 4]. Further discussions on the ABC condition can be found in [9] and [6].
Assumption 2.3.
$\sup_{z\in\pazocal{Z}}\ell(z,w_{*})<\infty$.
Assumption 2.3 is also used in [8]. Such an assumption can be satisfied when $\pazocal{Z}$ is bounded. See discussions in [8]. Note that by lemma C.6.3b, if $\ell(z,\cdot)$ is convex and assumption 2.1 holds, then the ‘noise’ $\|\delta m_{t}\|^{2}$ defined in (1.2) is bounded by a scalar multiple of $\|w_{t}-w_{*}\|^{2\gamma}+1$. This assumption will only be used to produce probability estimates for the convergence rate of SHB iterates.
2.2. Main Result 1: Almost Sure Convergence Rate
With assumption 2.1 and 2.2, in estimating $F(w_{t})-F_{*}$, one naturally arrives at a recursive inequality. In this regard, our key argument is a deterministic, weaker version of the Robbins-Siegmund theorem [13] and a lemma which deduce the rate.
Proposition 2.4.
Let $X_{t},Y_{t},Z_{t}$ be non-negative for all $t\in\mathbb{N}_{0}$ and $a_{t}>0$ be such that
  1. 1.
    $a_{t},Z_{t}\in\ell^{1}(\mathbb{N})$.
  2. 2.
    $Y_{t}\leq(1+a_{t-1})Y_{t-1}-X_{t-1}+Z_{t-1}$.
Then $Y_{t}\in\ell^{\infty}(\mathbb{N})$ and $X_{t}\in\ell^{1}(\mathbb{N})$.
Lemma 2.5.
Let $Y_{t},Z_{t}$ be a sequence of positive and non-negative random variables respectively, adapted to a filtration $\mathscr{F}_{t}$ and $c_{1},c_{2}\in\mathbb{R}_{+}$ be constants. Let $\alpha_{t}=\Theta(t^{-p})$ where $p\in(\frac{1}{1+\gamma},1)$ and $\gamma\in(0,1]$. If both of the following holds
  1. 1.
    For a fixed $s>0$, $\mathbb{E}[Y_{t+1}\,|\,\mathscr{F}_{s}]\leq\mathbb{E}[Y_{t}\,|\,\mathscr{F}_{s}](1+c_{1}\,\alpha_{t}^{1+\gamma})+\alpha_{t}^{1+\gamma}(\mathbb{E}[Z_{t}\,|\,\mathscr{F}_{s}]+c_{2})$ for all $t\geq s$.
  2. 2.
    $\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}[Y_{t}]<\infty$ and $\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}[Z_{t}]<\infty$.
then $\lim_{t\rightarrow\infty}t^{\min(1-p,p(1+\gamma)-1)-\epsilon}\,Y_{t}=0$.
Remark 2.6.
Compared to [9, lemma 19] and [13, theorem 1], we require $Y_{t}>0$ and item 2 where the expectation of the sum is required to be finite. The application of lemma 2.5 will be for convex cost $F$ and $Y_{t}=F(w_{t})-F_{*}$. Due to the requirement $Y_{t}>0$, instead of stating general convergence rate for $F(w_{t})-F_{*}$, we only make statements about the convergence rate of $F(w_{\tau\wedge t})-F_{*}$ where $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$.
Theorem 2.7.
Let $w_{t}$ be the iterates in (1.1). With the notation in table 2, if $\alpha_{t}=O(t^{-p})$ with $p\in(\frac{1}{1+\gamma},1)$ and both assumption 2.1 and 2.2 hold, then the SHB converges almost surely at rate
  1. 1.
    $\min_{0\leq s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ a.s.-$\mathbb{P}$.
  2. 2.
    If in addition, $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then
    1. (a)
      $\min_{0\leq s\leq t}F(w_{s})-F_{*}=o(t^{p-1})$ as $t\uparrow\infty$.
    2. (b)
      Let $\tau:=\inf\{t>0:F(w_{t})=F_{*}\}$. We have
      $F(w_{\tau\wedge t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})~~\text{where}~~r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0\end{cases}.$
Remark 2.8.
We include the ‘best case’ scenario (prefixed with $\min_{0\leq s\leq t}$) as part of our result. Best case scenario may be of less interest, but still provides information on how the rate compares with the ‘normal case’. In particular, it shows that the ‘optimal’ $t^{-1/2}$ rate for $\gamma=1$ is achievable if $p=\frac{1}{2}+\epsilon$. Further, in the application, one may stop the algorithm using a stopping time, in which case one may be interested in the ‘best case’ scenario.
Remark 2.9.
The slowdown factor $r_{\gamma}$ due to the smoothness parameter $\gamma$ appears only with momentum $\beta>0$ and $\gamma<1$. When $\gamma<1$, intuitively, the algorithm is slowed down as can be seen in the last term of (2.1). The fact that the slowdown factor $r_{\gamma}$ is uniform for all $\beta>0$ is interesting as empirical evidence showed that larger $\beta$ contributes to more slowdown [3]. There are also other works on non-acceleration of heavy ball [5]. We do not quite understand at the moment of these phenomena and how it ties to the slowdown factor $r_{\gamma}$. In addition, it is counter-intuitive since momentum is supposed to make convergence faster when the gradient is small. Examining (3.13), when there is momentum, the gradient $\|\nabla F(w_{t}+\frac{\beta}{1-\beta}(w_{t}-w_{t-1}))\|^{2}$ appears as an additive term in the upper estimate for $F(w_{t})-F_{*}$ and lemma C.4 suggests that for $\gamma<1$, this slows down the rate.
2.3. Main Result 2: Convergence Rate with High Probability
A key requirement to obtain fast asymptotic convergence is that $F(w_{t})-F_{*}$ is decreasing eventually. In other words, there exists $T>0$ such that $F(w_{t})$ is decreasing for all $t>T$. Such is the case in deterministic GD [4, lemma 2.28, eq. (14), algorithm 3.2]. In fact, smoothness of $F$ is sufficient (without convexity). Intuitively, in the stochastic GD, this would hold when $F$ is convex if the step size decay rate is controlled accordingly. However, even after taking expectation, this is difficult to achieve. Indeed, applying the expectation to (3.2) with [4, lemma 2.29] and proposition 3.1 yields an inequality of the form
$\mathbb{E}[F(w_{t+1})]-\mathbb{E}[F(w_{t})]$ $\leq-\alpha_{t}\bigg(\mathbb{E}\|\nabla F(w_{t})\|^{2}-\alpha_{t}^{\gamma}\left(A\,\mathbb{E}[F(w_{t})-F_{*}]+B\,\|\nabla F(w_{t})\|^{2}+C\right)\bigg)$
$\leq-\alpha_{t}\bigg(\min_{0\leq s\leq t}\mathbb{E}\|\nabla F(w_{s})\|^{2}-\alpha_{t}^{\gamma}\left(C+(A+2BL)\,\sup_{t\geq 0}\mathbb{E}[F(w_{t})-F_{*}]\right)\bigg).$
From [8, theorem 6], convex $F$ gives $\mathbb{E}\|\nabla F(w_{t})\|^{2}\leq\mathbb{E}[F(w_{t})-F_{*}]=O(t^{p-1})$ for $p>2/(2+\gamma)$. Since $p>1/(1+\gamma)$, we have $t^{p-1}>t^{p\gamma}$ which suggests that on average, $F(w_{t})$ decreases eventually.
We address this issue in our second result on SHB. Instead of looking at one-step iterates, we look at $F(w_{t})$ over a time period $t\in[\tau,T+1]$ in (4.35) by expressing
$F(w_{T+1})$ $\leq F(w_{\tau})+\sum_{t=\tau}^{T}\left(\beta\,\langle\nabla F(w_{t}),\,w_{t}-w_{t-1}\rangle+\frac{L}{1+\gamma}\,\|w_{t+1}-w_{t}\|^{1+\gamma}\right)$
$-\sum_{t=\tau}^{T}\left(\alpha_{t}\,\|\nabla F(w_{t})\|^{2}-\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle\right).$
The first summation on the right hand side decays quickly. The key is to control the last summation. Following the same idea as Lei, Shi, Guo in [8], we use concentration inequality of Bernstein and Azuma-Hoeffding to estimate the last term on the right hand side involving the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$. Effectively, the martingale difference term can be made smaller than $\alpha_{t}\,\|\nabla F(w_{t})\|^{2}$ so that the last term is effectively negative (more precisely, it is replaced with a quickly decaying term. See lemma 4.10). This suggests that with high probability, the ‘decreasing’ effect on $F(w_{t})$ (the effectiveness of SGD/SHB) is shown over a larger time step, which is a slowdown compared to determinsitic GD [4, theorem 3.4, lemma 2.28].
From the previous paragraph, applying martingale concentration inequalities on $\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ requires controlling the noise $\delta m_{t}$. By definition, $\delta m_{t}$ is comprised of $\nabla F(w_{t})$ and $\nabla\ell(Z_{t},w_{t})$, which by lemma C.4, is controlled by $\|w_{t}-w_{*}\|$. Following [8], in proposition 4.5, we show that up to time $T>0$, $\|w_{t}-w_{*}\|$ is at most $\log(T/\delta)$ with probability $1-\delta$. Since $\alpha_{t}$ decays with order $O(t^{-p})$, the martingale difference $\alpha_{t}\,\langle\nabla F(w_{t}),\,\delta m_{t}\rangle$ is of the order $\alpha_{t}$ up to a logarithmic factor.
Theorem 2.10.
Let $z_{t},w_{t},v_{t}$ be the SHB iterates as in (1.1) and (3.6) and $K_{0}(L,\gamma,\beta),\,K_{5}(L,\gamma,\beta)$ be positive constants defined in (4.1) and proposition 4.3. If $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, assumption 2.1 and assumption 2.3 both hold with $\gamma=1$, $\beta\in[0,1)$, and $\alpha_{t}$ satisfies
  1. 1.
    $\alpha_{0}\leq\min\left(1,\frac{1}{2\sqrt{K_{5}}},K_{0}\right)$.
  2. 2.
    $\alpha_{t+1}\leq\alpha_{t}$ for all $t\geq 0$.
  3. 3.
    $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{2},1)$.
then
$\mathbb{P}\left[F(w_{T+1})-F_{*}=O\left(T^{\max(p-1,-2p+1)}\left(\log\frac{T}{\delta}\right)^{2}\right)\right]\geq 1-\delta.$
Remark 2.11.
We remark that it is required to have $c_{1}t^{-p}\leq\alpha_{t}\leq c_{2}t^{-p}$ for all $t>0$, which is more stringent than the asymptotic $\Theta$-bound or $O$-bound since we need to estimate ratios of the form $\alpha_{t+s}/\alpha_{t}$ in (4.29) and $(\sum\alpha_{t})^{-1}$ in theorem 4.13. This result is consistent with theorem 2.7 and [9, theorem 13].
3. Almost Sure Convergence Rate
3.1. Stochastic Gradient Descent
Proposition 3.1.
If assumption 2.1, assumption 2.2 all hold and $\alpha_{t}\in\ell^{1+\gamma}(\mathbb{N})$ satisfies $\alpha_{t}\leq\min\{1,\frac{1}{\sqrt[\gamma]{LB}}\}$ for all $t\geq 0$, then the SGD algorithm (1.1) with $\beta=0$ satisfies
  1. 1.
    $\sup_{t\geq 0}\mathbb{E}[F(w_{t})-F_{*}]<\infty$ and $\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}\|\nabla F(w_{t})\|^{2}<\infty$.
  2. 2.
    If in addition $\alpha_{t}=O(t^{-p})$ where $p\in(\frac{1}{1+\gamma},1)$, then $\min_{1\leq s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ a.s.-$\mathbb{P}$.
  3. 3.
    If in addition $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then $\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]<\infty$.
Proof.
Due to $(\gamma,L)$-smoothness of $F$ and assumption 2.2, by proposition C.2,
$\mathbb{E}[F(w_{t})$ $-F_{*}-(F(w_{t-1})-F_{*})\,|\,\mathscr{F}_{t-1}]\leq\mathbb{E}[F(w_{t})-F(w_{t-1})\,|\,\mathscr{F}_{t-1}]$ (3.1)
$\leq\langle\nabla F(w_{t-1}),\,\mathbb{E}[w_{t}-w_{t-1}\,|\,\mathscr{F}_{t-1}]\rangle+\frac{L}{1+\gamma}\,\mathbb{E}[\|w_{t}-w_{t-1}\|^{1+\gamma}\,|\,\mathscr{F}_{t-1}]$
$\leq-\alpha_{t-1}\|\nabla F(w_{t-1})\|^{2}+\frac{L}{1+\gamma}\,\alpha_{t-1}^{1+\gamma}\,\left(A\,(F(w_{t-1})-F_{*})+B\,\|\nabla F(w_{t-1})\|^{1+\gamma}+C\right)$
By (2.4), we get
$\mathbb{E}[F(w_{t})-F_{*}$- $(F(w_{t-1})-F_{*})\,|\,\mathscr{F}_{t-1}]\leq-\alpha_{t-1}\|\nabla F(w_{t-1})\|^{2}$ (3.2)
$+\frac{L}{1+\gamma}\,\alpha_{t-1}^{1+\gamma}\,\left(A\,(F(w_{t-1})-F_{*})+2B\,\|\nabla F(w_{t-1})\|^{2}+2+C\right).$
If $Y_{t}=\mathbb{E}[F(w_{t})-F_{*}]$ and $LB\alpha_{t}^{\gamma}\leq 1$, then
$Y_{t}\leq\left(1+\frac{LA}{1+\gamma}\,\alpha_{t-1}^{1+\gamma}\right)\,Y_{t-1}-\left(\frac{\gamma}{1+\gamma}\right)\,\alpha_{t-1}\,\mathbb{E}\|\nabla F(w_{t-1})\|^{2}+\frac{L\,(C+2)}{1+\gamma}\,\alpha_{t-1}^{1+\gamma}.$ (3.3)
Applying proposition 2.4 yields the first claim. For the second claim, note that $\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty$ almost surely since its expectation is finite (due to the first claim). This implies
$\sum_{t=1}^{\infty}\left(\alpha_{t}\cdot\min_{1\leq s\leq t}\|\nabla F(w_{s})\|^{2}\right)\leq\sum_{t=1}^{\infty}\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty.$ (3.4)
Since in addition $\min_{1\leq s\leq t}\|\nabla F(w_{s})\|^{2}$ is monotonically decreasing, by [7, theorem 3.3.1], we have $t^{1-p}\,\min_{1\leq s\leq t}\|\nabla F(w_{s})\|^{2}\rightarrow 0$ almost surely.
For the third claim, let $W_{t}=\mathbb{E}\|w_{t}-w_{*}\|^{2}$. By the convexity of $F$, we obtain the estimate
$W_{t+1}$ $=W_{t}+2\,\mathbb{E}[\langle w_{t+1}-w_{t},\,w_{t}-w_{*}\rangle]+\mathbb{E}\|w_{t+1}-w_{t}\|^{2}$
$=W_{t}-2\alpha_{t}\,\mathbb{E}\langle\nabla F(w_{t}),\,w_{t}-w_{*}\rangle+\alpha_{t}^{2}\mathbb{E}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq W_{t}-2\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]+\alpha_{t}^{2}\left(A\,\mathbb{E}[F(w_{t})-F_{*}]+B\,\alpha_{t}\,\mathbb{E}\|\nabla F(w_{t})\|^{2}+C\right).$
Since $\alpha_{t}\in\ell^{2}(\mathbb{N})$, by proposition 3.1, the third term in the above estimate is summable. We may apply proposition 2.4 and conclude $\sum\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]<\infty$. ∎
Remark 3.2.
Note that the first two results in proposition 3.1 applies to non-convex cost function, comparable with [4, thm 5.12] by Garrigos and Gower for $\gamma=1$, which asserts that for a fixed end time $T\geq 1$, under a constant step size of the order $O(T^{-1/2})$, we have $\min_{0\leq t<T}\mathbb{E}[\|\nabla F(w_{t})\|^{2}]=O(T^{-1/2})$. More generally, our result is consistent with [9, theorem 6.2].
From proposition 3.1, we may conclude that $\mathbb{P}(\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty)=1$ and if $F$ is convex, $\mathbb{P}[\sum\alpha_{t}\,(F(w_{t})-F_{*})]=1$. Intuitively, this suggests that if $\sum\alpha_{t}=\infty$, then $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ almost surely (resp. for the convex case $F(w_{t})-F_{*}\rightarrow 0$ almost surely). Thanks to [12, theorem 2], such intuition holds in our case.
Corollary 3.3.
If all assumptions in proposition 3.1 holds and $\alpha_{t}\notin\ell^{1}(\mathbb{N})$, then
  1. 1.
    $\|\nabla F(w_{t})\|\rightarrow 0$ almost surely.
  2. 2.
    If $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then
    1. (a)
      $F(w_{t})-F_{*}\rightarrow 0$ almost surely.
    2. (b)
      Let $\tau$ be defined as in theorem 2.7. If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then almost surely,
      $F(w_{\tau\wedge t})-F_{*}=o(t^{\max(p-1,1-(1+\gamma)p)+\epsilon}).$
Proof.
Following the same argument as in [12, theorem 2], we get
$\bigg|\|\nabla F(w_{t})\|-\|\nabla F(w_{s})\|\bigg|^{1/\gamma}$ $\leq\|\nabla F(w_{t})-\nabla F(w_{s})\|^{1/\gamma}\leq L\sum_{i=s}^{t}\alpha_{i}\,\|\nabla F(w_{i})\|+L\,\left\|\sum_{i=s}^{t}\alpha_{i}\,\delta m_{i}\right\|.$
Since proposition 3.1 asserts $\sum\alpha_{t}\,\|\nabla F(w_{t})\|^{2}<\infty$ almost surely, by lemma C.3, it suffices to show that the second term on the right hand side has finite $L^{2}$ norm. Indeed, if this is the case, then $\sum_{i=s}^{t}\alpha_{i}\delta m_{i}<\infty$ almost surely. By proposition C.7 and assumption 2.2,
$\mathbb{E}\left\|\sum_{i=s}^{t}\alpha_{i}\,\delta m_{i}\right\|^{2}=\sum_{i=s}^{t}\alpha_{i}^{2}\,\mathbb{E}\|\delta m_{i}\|^{2}\leq\sum_{i=s}^{t}\alpha_{i}^{2}\left(A\,\mathbb{E}[F(w_{i})-F_{*}]+B\,\mathbb{E}\|\nabla F(w_{i})\|^{2}+C\right)$
The term on the right hand side is finite due to the first conclusion in proposition 3.1 and that $\alpha_{t}\in\ell^{1+\gamma}\subset\ell^{2}$. Therefore, the first claim follows.
Part 2.a. For the convex case, we know that $F$ has no saddle points and all minima are global. Therefore, the set of critical points $\pazocal{C}=\{w\in\pazocal{W}:\|\nabla F(w)\|^{2}=0\}$ is the set of global minima. Therefore, $\|\nabla F(w_{t})\|^{2}\rightarrow 0$ implies $w_{t}\rightarrow w$ where $w\in\pazocal{C}$. By continuity of $F$, we get $F(w_{t})\rightarrow F(w)=F_{*}$.
Part 2.b. For a realization $\omega\in\Omega$, if $\tau(\omega)<\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F_{*}$ for $t\geq\tau(\omega)$. The claim then holds trivially. If $\tau(\omega)=\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F(w_{t}(\omega))$ for all $t>0$. Defining $Y_{t}=F(w_{t})-F_{*}$, applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.2), dropping the $\alpha_{t}\,\|\nabla F(w_{t})\|^{2}$ term, and applying lemma C.4 yields an inequality of the form
$\mathbb{E}[Y_{t+1}\,|\,\mathscr{F}_{s}]$ $\leq\left(1+c_{0}\,\alpha_{t}^{1+\gamma}\right)\mathbb{E}[Y_{t}\,|\,\mathscr{F}_{s}]+c_{0}\,\alpha_{t}^{1+\gamma}\bigg(c_{1}\,\mathbb{E}[Y_{t}\,|\,\mathscr{F}_{s}]+c_{2}\bigg)$
$=\left(1+(c_{0}+c_{0}c_{1})\alpha_{t}^{1+\gamma}\right)\,\mathbb{E}[Y_{t}\,|\,\mathscr{F}_{s}]+c_{0}c_{2}\,\alpha_{t}^{1+\gamma}$
where $c_{0}=\frac{L}{1+\gamma}$, $c_{1}=A+4BL$, and $c_{2}=C+2$. By proposition 3.1, all conditions in lemma 2.5 are satisfied. Therefore, the claim holds. ∎
3.2. Stochastic Heavy Ball
The algorithm for SHB is given in (1.1). From [9, eq. (17) and (18)], by defining
$\begin{cases}v_{t}=w_{t}-w_{t-1}~,~z_{t}=w_{t}+\frac{\beta}{1-\beta}\,v_{t}~~,~~t\geq 1.\\ w_{1}=w_{0}.\end{cases}$ (3.5)
the SHB can be rewritten as one-step iterates
$\begin{cases}v_{t+1}=\beta\,v_{t}-\alpha_{t}\,\nabla\ell(Z_{t},w_{t})&t\geq 1.\\ z_{t+1}=z_{t}-\displaystyle\frac{\alpha_{t}}{1-\beta}\,\nabla\ell(Z_{t},w_{t})&t\geq 1.\\ z_{1}=w_{1}~,~v_{1}=0.\end{cases}$ (3.6)
Proposition 3.4.
If assumption 2.1 and assumption 2.2 holds, then there exists a constant $b_{0}>0$ such that if $\alpha_{t}\in\ell^{1+\gamma}$ and $\alpha_{t}\leq\min(1,b_{0})$, then the SHB algorithm (3.6) corresponding to (1.1) satisfies
  1. 1.
    $\sup_{t\geq 0}\mathbb{E}[F(z_{t})-F_{*}+\|v_{t}\|^{2}]<\infty$ and $\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}\|\nabla F(z_{t})\|^{2}<\infty$.
  2. 2.
    $\sum_{t=1}^{\infty}\mathbb{E}\|v_{t}\|^{2}<\infty$ a.s.-$\mathbb{P}$.
  3. 3.
    If $\alpha_{t}=O(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then $\min_{1\leq s\leq t}\|\nabla F(w_{s})\|^{2}=o(t^{p-1})$ a.s.-$\mathbb{P}$.
  4. 4.
    If $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then $\sum_{t=1}^{\infty}\alpha_{t}\left(\mathbb{E}[F(w_{t})-F_{*}]+\mathbb{E}[F(z_{t})-F_{*}]\right)<\infty$.
Proof.
Proof of claim 1. Following [9, eq. (21)-(22)] in the proof of [9, theorem 8] and using the simple estimate $a^{1+\gamma}\leq 2a^{2}+2$ as in (2.4), we have
$\|\nabla F(w_{t})\|^{2}$ $\leq 2\|\nabla F(z_{t})\|^{2}+c_{1}\,\|v_{t}\|^{2}$ (3.7)
$F(w_{t})-F_{*}$ $\leq F(z_{t})-F_{*}+\frac{\|\nabla F(z_{t})\|^{2}}{2}+\left\{c_{2}+c_{3}\right\}\,\|v_{t}\|^{2}+c_{4}$ (3.8)
where
$c_{1}:=\frac{2L^{2}\beta^{2}}{(1-\beta)^{2}}~~,~~c_{2}:=\frac{\beta^{2}}{2(1-\beta)^{2}}~~,~~c_{3}:=\frac{L\beta^{1+\gamma}}{(1+\gamma)(1-\beta)^{1+\gamma}}~~,~~c_{4}:=2(c_{3}+1).$
Observe that the left hand sides of (3.7) and (3.8) makeup the bounding terms in assumption 2.2. This implies that the upper bound of $\mathbb{E}[\nabla\ell(Z_{t},w_{t})\,|\,\mathscr{F}_{t}]$ can be expressed entirely in terms of SHB variables $F(z_{t})-F_{*}$, $\|\nabla F(z_{t})\|^{2}$, and $\|v_{t}\|^{2}$. As a consequence, we may estimate the iterates of the SHB variables as follows
$\mathbb{E}[\|v_{t+1}\|^{2}\,|\,\mathscr{F}_{t}]$ $\leq\beta^{2}\,(1+\varepsilon_{1})\,\|v_{t}\|^{2}+\frac{1}{\varepsilon_{1}}\,\alpha_{t}^{2}\,\|\nabla F(w_{t})\|^{2}$ (3.9)
$+\alpha_{t}^{2}\left(A\,[F(w_{t})-F_{*}]+B\,\|\nabla F(w_{t})\|^{2}+C\right)$
$\leq\left(\beta^{2}\,(1+\varepsilon_{1})+\left\{\left(\frac{1}{\varepsilon_{1}}+B\right)c_{1}+A\,c_{2}+A\,c_{3}\right\}\,\alpha_{t}^{2}\right)\|v_{t}\|^{2}$
$+\left(\frac{2}{\varepsilon_{1}}+2B+\frac{A}{2}\right)\,\alpha_{t}^{2}\,\|\nabla F(z_{t})\|^{2}+\alpha_{t}^{2}\,\bigg(C+A\,c_{4}\bigg)$
$+A\,\alpha_{t}^{2}\,[F(z_{t})-F_{*}]$
where $\varepsilon_{1}$ is an arbitrary constant. In addition, following [9, eq.(20)], we get
$\mathbb{E}[F(z_{t+1})\,|\,\mathscr{F}_{t}]$ $\leq F(z_{t})-\left(\frac{1}{1-\beta}-\frac{\alpha_{t}\,c_{5}}{\varepsilon_{2}}\right)\alpha_{t}\,\|\nabla F(z_{t})\|^{2}+\varepsilon_{2}\|v_{t}\|^{2}$ (3.10)
$+c_{6}\,\alpha_{t}^{1+\gamma}\,(A[F(w_{t})-F_{*}]+B\|\nabla F(w_{t})\|^{2}+C)$
$\leq F(z_{t})-\left(\frac{1}{1-\beta}-\frac{\alpha_{t}\,c_{5}}{\varepsilon_{2}}-\left(B\,c_{6}+\frac{A}{2}\right)\,\alpha_{t}^{\gamma}\right)\alpha_{t}\,\|\nabla F(z_{t})\|^{2}$
$+\bigg(\varepsilon_{2}+(c_{1}\,c_{6}\,B+A\,c_{2}+A\,c_{3})\,\alpha_{t}^{1+\gamma}\bigg)\,\|v_{t}\|^{2}$
$+A\,c_{6}\,\alpha_{t}^{1+\gamma}\,[F(z_{t})-F_{*}]+(C+A\,c_{4})\,\alpha_{t}^{1+\gamma}$
where $\varepsilon_{2}$ is an arbitrary constant and
$c_{5}:=\frac{c_{1}(L,\beta)}{8(1-\beta)}~,~c_{6}:=\frac{c_{3}(L,\beta,\gamma)}{\beta^{1+\gamma}}$
Since $\beta<1$, we take $\varepsilon_{1},\varepsilon_{2}\in\mathbb{R}_{+}$ such that $\beta^{2}(1+\varepsilon_{1})+\varepsilon_{2}\leq 1$, e.g. set $\varepsilon_{1}=\frac{1-\beta^{2}}{2\beta^{2}}$ and $\varepsilon_{2}=\frac{1-\beta^{2}}{2}$. Now, adding (3.9) and (3.10) yields that there exists positive constants $b_{0},b_{1},b_{2},b_{3}\in\mathbb{R}_{+}$ such that if $\alpha_{t}\leq\min(1,b_{0})$, then
$\mathbb{E}[F(z_{t+1})-F_{*}+\|v_{t+1}\|^{2}\,|\,\mathscr{F}_{t}]$ $\leq(1+b_{1}\,\alpha_{t}^{1+\gamma})(F(z_{t})-F_{*}+\|v_{t}\|^{2})$ (3.11)
$-b_{2}\,\alpha_{t}\,\|\nabla F(z_{t})\|^{2}+b_{3}\,\alpha_{t}^{1+\gamma}.$
Applying the expectation on both sides and proposition 2.4 yields the first claim.
Proof of claim 2 and 3. Observe that the iterates of $v_{t}$ satisfies
$\mathbb{E}\|v_{t+1}\|^{2}$ $=\beta^{2}\,\mathbb{E}\|v_{t}\|^{2}-2\beta\,\alpha_{t}\,\mathbb{E}[\langle\nabla F(w_{t}),\,v_{t}\rangle]+\alpha_{t}^{2}\,\mathbb{E}\|\nabla\ell(Z,w_{t})\|^{2}$ (3.12)
$\leq\mathbb{E}\|v_{t}\|^{2}-2\beta\,\alpha_{t}\,\mathbb{E}[\langle\nabla F(w_{t}),\,v_{t}\rangle]+\alpha_{t}^{2}\,\mathbb{E}\|\nabla\ell(Z,w_{t})\|^{2}.$
Using assumption 2.2, the first claim, along with $\ell^{1+\gamma}\subset\ell^{2}$, (3.7), and (3.8) yields that the last term in (3.12) is summable:
$\sum_{t=1}^{\infty}\alpha_{t}^{2}\,\bigg(\mathbb{E}\|\nabla F(z_{t})\|^{2}+\sup_{t\geq 0}\,\mathbb{E}\left[F(z_{t})-F_{*}+\|v_{t}\|^{2}\right]+1\bigg)<\infty.$
Therefore, by proposition 2.4, the middle term in (3.12) is also summable: $\sum\alpha_{t}\mathbb{E}[\langle\nabla F(w_{t}),\,v_{t}\rangle]<\infty$. Now we have that $\mathbb{E}\|v_{t+1}\|^{2}-\beta^{2}\mathbb{E}\|v_{t}\|^{2}$ corresponds to summable terms. Since $v_{1}=0$, we know $\sum_{t=1}^{\infty}\|v_{t}\|^{2}=\sum_{t=1}^{\infty}\|v_{t+1}\|^{2}$. Therefore,
$(1-\beta^{2})\sum_{t=1}^{\infty}\mathbb{E}\|v_{t}\|^{2}=-2\beta\sum_{t=1}^{\infty}\alpha_{t}\,\mathbb{E}[\langle\nabla F(w_{t}),\,v_{t}\rangle]+\sum_{t=1}^{\infty}\alpha_{t}^{2}\,\mathbb{E}\|\nabla\ell(Z_{t},w_{t})\|^{2}<\infty.$
This proves $\sum\mathbb{E}\|v_{t}\|^{2}<\infty$, which implies $\sum_{t=1}^{\infty}\|v_{t}\|^{2}<\infty$ a.s.-$\mathbb{P}$. The third claim follows from (3.7) using similar argument as in the proof of the second result in proposition 3.1.
Proof of claim 4. Define $U_{t}=\mathbb{E}\|z_{t}-w_{*}\|^{2}$ and $K_{ABC}=C+(A+2BL)\sup_{t\geq 0}\mathbb{E}[F(w_{t})-F_{*}]$. By the convexity of $F$,
$U_{t+1}$ $=U_{t}-2\alpha_{t}\,\mathbb{E}[\langle\nabla F(w_{t}),\,z_{t}-w_{*}\rangle]+\alpha_{t}^{2}\,\mathbb{E}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$=U_{t}-2\alpha_{t}\left(\mathbb{E}[\langle\nabla F(w_{t}),\,w_{t}-w_{*}\rangle]+\frac{\beta}{1-\beta}\mathbb{E}[\langle\nabla F(w_{t}),\,v_{t}\rangle]\right)+\alpha_{t}^{2}\mathbb{E}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq U_{t}-2\,\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]+\left(\frac{2\beta}{1-\beta}\alpha_{t}(F(w_{t-1})-F(w_{t}))+\alpha_{t}^{2}\,K_{ABC}\right)$
$\leq U_{t}-2\,\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]+\left(\frac{2\beta}{1-\beta}\bigg(\alpha_{t-1}\mathbb{E}[F(w_{t-1})]-\alpha_{t}\mathbb{E}[F(w_{t})]\bigg)+\alpha_{t}^{2}\,K_{ABC}\right).$
Note that we have used convexity of $F$ and that $\alpha_{t}$ is decreasing in the second inequality. Note that terms in the parentheses are summable. Applying proposition 2.4 yields that $\sum\alpha_{t}\,\mathbb{E}[F(w_{t})-F_{*}]<\infty$ for which the fourth claim follows. Lastly, $(\gamma,L)$-smoothness gives
$F(z_{t})-F_{*}$ $\leq F(w_{t})-F_{*}+\langle\nabla F(w_{t}),\,z_{t}-w_{t}\rangle+\frac{L}{1+\gamma}\|z_{t}-w_{t}\|^{1+\gamma}$
$\leq F(w_{t})-F_{*}+\frac{\|\nabla F(w_{t})\|^{2}}{2}+\frac{\beta^{2}}{2(1-\beta)^{2}}\|v_{t}\|^{2}+\frac{L}{1+\gamma}\left(\frac{\beta}{1-\beta}\right)^{1+\gamma}\|v_{t}\|^{1+\gamma}.$
Multiplying both sides by $\alpha_{t}$ yields that the right hand side is summable. ∎
Remark 3.5.
Proposition 3.4 shows that the SHB converges at rate similar to that of SGD. In particular, the SHB parameter $\beta$ in (1.1) does not impact the almost sure convergence rate, consistent with Liu and Yuan’s result [9, theorem 8].
Corollary 3.6.
If all the assumptions in proposition 3.4 hold and $\alpha_{t}\notin\ell^{1}(\mathbb{N})$, then
  1. 1.
    $\|\nabla F(w_{t})\|\rightarrow 0$ almost surely.
  2. 2.
    If $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then
    1. (a)
      $F(w_{t})-F_{*}\rightarrow 0$ almost surely.
    2. (b)
      Let $\tau$ be defined as in theorem 2.7. If $\alpha_{t}=\Theta(t^{-p})$ for $p\in(\frac{1}{1+\gamma},1)$, then almost surely
      $F(w_{\tau\wedge t})-F_{*}=o(t^{r_{\gamma}\max(p-1,1-(1+\gamma)p)+\epsilon})~~\text{where}~~r_{\gamma}=\begin{cases}\frac{2\gamma}{1+\gamma}&\beta\in(0,1)\\ 1&\beta=0\end{cases}.$
Proof.
The same argument as in corollary 3.3 can be used. For a realization $\omega\in\Omega$, if $\tau(\omega)<\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F_{*}$ for all $t\geq\tau(\omega)$ and the claim holds trivially. Otherwise, if $\tau(\omega)=\infty$, then $F(w_{\tau(\omega)\wedge t}(\omega))=F(w_{t}(\omega))$ so that the argument in the next paragraph can be used.
Both claims follow from (3.7) and (3.8) along with the first and second conclusion in proposition 3.4. Let $Y_{t}=F(z_{t})-F_{*}+\|v_{t}\|^{2}$. Applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{s}]$ on (3.11) and dropping the $\alpha_{t}\|\nabla F(z_{t})\|^{2}$ term yields
$\mathbb{E}[Y_{t+1}\,|\,\mathscr{F}_{s}]\leq(1+b_{1}\,\alpha_{t}^{1+\gamma})\,\mathbb{E}[Y_{t}\,|\,\mathscr{F}_{s}]+b_{3}\,\alpha_{t}^{1+\gamma}.$
Proposition 3.4 shows that all conditions in lemma 2.5 holds. Estimating $F(w_{t})-F_{*}$ using $(\gamma,L)$-smoothness and applying lemma C.4 yields
$F(w_{t})-F_{*}$ $\leq F(z_{t})-F_{*}+\frac{\beta}{1-\beta}\langle\nabla F(z_{t}),\,v_{t}\rangle+\frac{L}{1+\gamma}\left(\frac{\beta}{1-\beta}\|v_{t}\|\right)^{1+\gamma}$ (3.13)
$\leq F(z_{t})-F_{*}+\frac{\beta}{1-\beta}\left(\frac{\|\nabla F(z_{t})\|^{2}+\|v_{t}\|^{2}}{2}\right)+\frac{L}{1+\gamma}\left(\frac{\beta}{1-\beta}\|v_{t}\|\right)^{1+\gamma}$
$\leq F(z_{t})-F_{*}+\frac{\beta}{1-\beta}\left(\frac{L^{1/\gamma}}{1+\gamma}\cdot(F(z_{t})-F_{*})^{\frac{2\gamma}{1+\gamma}}+\frac{\|v_{t}\|^{2}}{2}\right)+\frac{L}{1+\gamma}\left(\frac{\beta}{1-\beta}\|v_{t}\|\right)^{1+\gamma}.$
Applying lemma 2.5 yields $t^{r}(F(w_{t})-F_{*})\rightarrow 0$ if
$r=\begin{cases}\min(\frac{1+\gamma}{2},\frac{2\gamma}{1+\gamma})\cdot\min(1-p,p(1+\gamma)-1)&\beta\in(0,1)\\ \min(1-p,p(1+\gamma)-1)&\beta=0.\end{cases}$
This completes the proof. ∎
4. Convergence Rate with High Probability for Convex Objective
We will assume throughout this section that $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, which by proposition C.1, makes $F$ also convex. It will be convenient to define once and for all the following constants and notations. Let $a_{1}$ and $a_{2}$ be the constants defined in lemma C.6. Also, define the following
$\begin{aligned} k_{0}&:=\frac{\beta}{1-\beta}~~,~~k_{1}:=\frac{a_{1}(L,\gamma)}{(1-\beta)^{2}}~~,~~k_{2}:=\frac{a_{2}(L,\gamma)}{(1-\beta)^{2}}~~,~~k_{3}:=\frac{2}{1-\beta}\cdot\sup_{z\in\pazocal{Z}}\ell(z,w_{*}).\\ k_{4}&:=k_{0}+\beta\,k_{0}^{2}~~,~~k_{5}:=k_{2}+k_{3}+a_{2}\,k_{0}^{2}.\\ K_{0}&:=\min\left(1,\frac{1}{(a_{1}\,k_{0}^{2}+a_{1}\,k_{0}+a_{1}+k_{1})(1-\beta)}\right).\\ \Delta z_{t}&=\|z_{t}-w_{*}\|^{2}-\|z_{t-1}-w_{*}\|^{2}~~,~~\Delta v_{t}=\|v_{t}\|^{2}-\|v_{t-1}\|^{2}.\end{aligned}$ (4.1)
4.1. Distance from a Minimizer with High Probability
Lemma 4.1.
Let $z_{t},w_{t},v_{t}$ be the SHB iterates as in (1.1) and (3.6). With the notation in (4.1), if $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, assumption 2.1 and assumption 2.3 both hold, and $\alpha_{t}$ satisfies
  1. 1.
    $\alpha_{0}\leq K_{0}$.
  2. 2.
    $\alpha_{t+1}\leq\alpha_{t}$ for all $t\geq 0$.
  3. 3.
    $\alpha_{t}\in\ell^{2}(\mathbb{N})$.
then there exists constants $K_{1}(L,\beta,\gamma,w_{0})>0$ and $K_{2}(L,\beta,\gamma,w_{0})>0$ such that for all $t>0$,
  1. 1.
    $\max\left\{\|w_{t}-w_{*}\|^{2},\|v_{t}\|^{2},\|z_{t}\|^{2}\right\}\leq K_{1}\,\sum_{s=0}^{t-1}\alpha_{s}$ a.s.-$\mathbb{P}$.
  2. 2.
    $\sum_{s=1}^{t-1}\alpha_{s}^{2}\,\ell(Z_{s},w_{s})\leq K_{2}$ a.s.-$\mathbb{P}$.
Proof.
Consider (3.6) for the iterates of $v_{t}$. Take the squared-norm. Applying lemma C.6.3a yields
$\|v_{t+1}\|^{2}-\|v_{t}\|^{2}$ $\leq\|v_{t+1}\|^{2}-\beta^{2}\,\|v_{t}\|^{2}$ (4.2)
$=-2\,\alpha_{t}\,\beta\,\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle+\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq-2\,\alpha_{t}\,\beta\,\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle+a_{1}\,\alpha_{t}^{2}\,\ell(Z_{t},w_{t})+a_{2}\,\alpha_{t}^{2}$
Using the notation in (4.1), convexity and lemma C.6.3a yields
$\Delta z_{t+1}$ $=-2\,\alpha_{t}\,\langle\nabla\ell(Z_{t},w_{t}),\,\frac{\beta v_{t}+w_{t}-w_{*}}{1-\beta}\rangle+\frac{1}{(1-\beta)^{2}}\,\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}$ (4.3)
$\leq-2\,\alpha_{t}\,\langle\nabla\ell(Z_{t},w_{t}),\,\frac{\beta v_{t}}{1-\beta}\rangle+\frac{2}{1-\beta}\,\alpha_{t}\bigg(\ell(Z_{t},w_{*})-\ell(Z_{t},w_{t})\bigg)+k_{1}\,\alpha_{t}^{2}\,\ell(Z_{t},w_{t})+k_{2}\,\alpha_{t}^{2}$
$\leq-2\,\alpha_{t}\,\langle\nabla\ell(Z_{t},w_{t}),\,\frac{\beta v_{t}}{1-\beta}\rangle+\alpha_{t}\,(k_{3}+k_{2}\,\alpha_{t})-\alpha_{t}\,\ell(Z_{t},w_{t})\bigg(\frac{2}{1-\beta}-\alpha_{t}\,k_{1}\bigg).$
Now, let $A_{t}=\sum_{s=1}^{t-1}\alpha_{s}$. Note that $v_{1}=0$ and $z_{1}=w_{0}$. Multiplying (4.2) by $k_{0}^{2}$, adding (4.3), and summing from $s=1$ to s=t1s=t-1 yields
$\|z_{t}-w_{*}\|^{2}+k_{0}^{2}\,\|v_{t}\|^{2}$ $\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}$ (4.4)
$-2\beta\left(\frac{1}{1-\beta}+k_{0}^{2}\right)\sum_{s=1}^{t-1}\alpha_{s}\,\langle\nabla\ell(Z_{s},w_{s}),\,v_{s}\rangle$
$-\sum_{s=1}^{t-1}\alpha_{s}\,\ell(Z_{s},w_{s})\bigg(\frac{2}{1-\beta}-\alpha_{s}\,(k_{1}+a_{1}k_{0}^{2})\bigg).$
Let $X_{t}$ be the last term in (4.4) and $Y_{t}=2k_{0}\langle w_{0}-w_{*},\,v_{t}\rangle$. Using (4.4) and $z_{1}=w_{0}$, we get the estimate
$\|w_{t}$ $-w_{*}\|^{2}=\|z_{t}-w_{*}\|^{2}+k_{0}^{2}\,\|v_{t}\|^{2}-2k_{0}\,\langle z_{t}-w_{*},\,v_{t}\rangle$ (4.5)
$=\|z_{t}-w_{*}\|^{2}+k_{0}^{2}\,\|v_{t}\|^{2}-2k_{0}\langle z_{1}+w_{*},\,v_{t}\rangle-2k_{0}\sum_{s=1}^{t-1}\langle z_{s+1}-z_{s},\,v_{s}\rangle$
$=\|z_{t}-w_{*}\|^{2}+k_{0}^{2}\,\|v_{t}\|^{2}-2k_{0}\langle w_{0}-w_{*},\,v_{t}\rangle+\frac{2k_{0}}{1-\beta}\sum_{s=1}^{t-1}\alpha_{s}\,\langle\nabla\ell(Z_{s},w_{s}),\,v_{s}\rangle$
$\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}+X_{t}-Y_{t}+2\,\beta\,k_{0}\sum_{s=1}^{t-1}\alpha_{s}\,\langle\nabla\ell(Z_{s},w_{s}),\,v_{s}\rangle.$
Now, recalling $v_{1}=0$ and rearranging (4.2) and summing from $s=1$ to s=t1s=t-1 yields
$2\beta\,\sum_{s=1}^{t-1}\alpha_{s}\langle\nabla\ell(Z_{s},w_{s}),\,v_{s}\rangle$ $\leq-\|v_{t}\|^{2}+a_{1}\,\sum_{s=1}^{t-1}\alpha_{s}^{2}\,\ell(Z_{s},w_{s})+a_{2}\sum_{s=1}^{t-1}\alpha_{s}^{2}.$
Combining this with (4.5) yields that the middle term $a_{1}\sum\alpha_{s}^{2}\ell(Z_{s},w_{s})$ can be collected with $X_{t}$ so that
$\|w_{t}-w_{*}\|^{2}$ $\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}-Y_{t}-2k_{0}\|v_{t}\|^{2}-k_{0}a_{2}\sum_{s=1}^{t-1}\alpha_{s}^{2}$ (4.6)
$-\sum_{s=1}^{t-1}\alpha_{s}\,\ell(Z_{s},w_{s})\bigg(\frac{2}{1-\beta}-\alpha_{s}\,(k_{1}+a_{1}k_{0}^{2}+a_{1}k_{0})\bigg).$
Requiring $\alpha_{t}$ to be small enough, we may drop the last two terms on the right hand side of (4.6). Now, using the inequality $2|\langle w_{0}-w_{*},\,v_{t}\rangle|\leq\varepsilon\|w_{0}-w_{*}\|^{2}+\frac{1}{\varepsilon}\|v_{t}\|^{2}$ and setting $\varepsilon=1/2$ yields
$\|w_{t}-w_{*}\|^{2}$ $\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}-Y_{t}-2k_{0}\|v_{t}\|^{2}$ (4.7)
$\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}+2k_{0}|\langle w_{0}-w_{*},\,v_{t}\rangle|-2k_{0}\|v_{t}\|^{2}$
$\leq\|w_{0}-w_{*}\|^{2}+k_{5}\,A_{t}+\frac{k_{0}}{2}\|w_{0}-w_{*}\|^{2}.$
Since we define $\alpha_{0}>0$, we have $\|w_{t}-w_{*}\|^{2}\lesssim\sum_{s=0}^{t-1}\alpha_{s}$ as desired. Since $v_{t}=w_{t}-w_{t-1}$, we know that $\|v_{t}\|^{2}$ has the same order as $\|w_{t}-w_{*}\|^{2}$. Similarly, since $z_{t}=w_{t}+k_{0}\,v_{t}$, it follows that $\|z_{t}\|^{2}$ also has the same order as $\|w_{t}-w_{*}\|^{2}$.
Now, define $\Delta v_{t+1}^{\beta}=\beta^{2}\|v_{t}\|^{2}-\|v_{t+1}\|^{2}$. Observe that
$\Delta z_{t+1}-\Delta v_{t+1}^{\beta}$ $\leq(k_{3}+k_{2}\,\alpha_{t})\alpha_{t}-\alpha_{t}\,\ell(Z_{t},w_{t})\left(\frac{2}{1-\beta}-\alpha_{t}\,k_{1}\right)-\alpha_{t}^{2}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq k_{5}\,\alpha_{t}-\alpha_{t}\,\ell(Z_{t},w_{t})\left(\frac{2}{1-\beta}-\alpha_{t}\,k_{1}\right)+\alpha_{t}^{2}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq(k_{5}+a_{2}\alpha_{t})\,\alpha_{t}-\alpha_{t}\,\ell(Z_{t},w_{t})\left(\frac{2}{1-\beta}-\alpha_{t}\,(k_{1}+a_{1})\right).$
If $\alpha_{t}$ is small enough and is decreasing, rearranging the above by moving the last term to the left hand side and multiplying by $\alpha_{t}$ yields
$\left(\frac{1}{1-\beta}\right)\alpha_{t}^{2}\,\ell(Z_{t},w_{t})$ $\leq(k_{5}+a_{2}\,\alpha_{t})\,\alpha_{t}^{2}-\Delta z_{t+1}+\Delta v_{t+1}^{\beta}$ (4.8)
$=(k_{3}+a_{2}\,\alpha_{t})\,\alpha_{t}^{2}-\alpha_{t}\Delta z_{t+1}+\alpha_{t}\beta^{2}\|v_{t}\|^{2}-\alpha_{t}\|v_{t+1}\|^{2}$
$\leq(k_{3}+a_{2})\,\alpha_{t}^{2}+\alpha_{t}\|z_{t}\|^{2}-\alpha_{t+1}\|z_{t+1}\|^{2}+\alpha_{t}\|v_{t}\|^{2}-\alpha_{t+1}\|v_{t+1}\|^{2}.$
Since $\alpha_{t}\in\ell^{2}$, summing the above and making use of the first claim completes the proof. ∎
Remark 4.2.
Note that the $\beta$-parameter does not affect the long term fluctuation of the iterates as can be seen from the conclusion of lemma 4.1 and from setting $\beta=0$ directly in estimates (4.7) and (4.8). As such, lemma 4.1 holds if $w_{t}$ is the SGD iterates.
In the sequel, we will require estimating constants of the form $\alpha_{t}\sum_{s=1}^{t}\alpha_{s}$. Under the assumption $\alpha_{t}\in\ell^{1+\gamma}(\mathbb{N})$ for $\gamma\in(0,1]$, we have $\alpha_{t}\sum_{s=0}^{t-1}\alpha_{s}\leq\alpha_{t}^{1-\gamma}\sum_{s=0}^{t-1}\alpha_{s}^{1+\gamma}<\infty$. Moreover, such a sum converges to zero for $\gamma\in(0,1)$. If $\gamma=1$, $\alpha_{0}=K_{0}$, $\alpha_{t}\leq K_{0}\,t^{-p}$ for $t>0$, then
$\alpha_{t}\sum_{s=0}^{t-1}\alpha_{s}\leq K_{0}^{2}\,t^{-p}\left(1+\sum_{s=1}^{t-1}s^{-p}\right)\lesssim K_{0}^{2}\,\max(t^{-p},(t-1)^{-2p+1})\leq K_{0}^{2}\leq K_{0}~~~,~~~\forall t>0~~~.$ (4.9)
In addition, from here onwards, we will estimate probabilities using Azuma-Hoeffding and Bernstein’s inequality following the work by Lei, Shi, Guo in [8].
Proposition 4.3.
With the notation in lemma 4.1, if all assumptions in lemma 4.1 holds and in addition $\alpha_{t}\leq K_{0}t^{-p}$ for $p\in(\frac{1}{2},1)$, then there exists positive constants $K_{3},K_{4},K_{5}\in\mathbb{R}_{+}$ depending only on $L,\gamma,\beta$ such that for any $t>0$ and $\delta\in(0,1)$,
$\mathbb{P}\left(\|z_{t}-w_{*}\|^{2}+\left(\frac{\beta}{1-\beta}\right)^{2}\,\|v_{t}\|^{2}\leq K_{3}+K_{4}\log\frac{1}{\delta}+K_{5}\sum_{s=1}^{t-1}\alpha_{s}^{2}\,\|w_{s}-w_{*}\|^{2}\right)\geq 1-\delta.$
Proof.
We use the notation of constants in (4.1). Using the convexity of $F$, summing (4.2) and (4.3) yields
$\Delta z_{t+1}+k_{0}^{2}\,\Delta v_{t+1}$ $\leq-2\,\alpha_{t}\,\beta\left(\frac{1}{1-\beta}+k_{0}^{2}\right)\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle-\frac{2\alpha_{t}}{1-\beta}\langle\nabla\ell(Z_{t},w_{t}),\,w_{t}-w_{*}\rangle+\varphi_{t}$ (4.10)
$=-2\,\alpha_{t}\,k_{4}\,\langle\nabla F(w_{t}),\,w_{t}-w_{t-1}\rangle-\frac{2\alpha_{t}}{1-\beta}\,\langle\nabla F(w_{t}),\,w_{t}-w_{*}\rangle+\frac{2\xi_{t}}{1-\beta}+\varphi_{t}$
$\leq 2\,k_{4}\,\alpha_{t}\big(F(w_{t-1})-F(w_{t})\big)+\frac{2\alpha_{t}}{1-\beta}\big(F(w_{*})-F(w_{t})\big)+\frac{2\xi_{t}}{1-\beta}+\varphi_{t}$
$\leq 2\,k_{4}\,\big(\alpha_{t-1}F(w_{t-1})-\alpha_{t}F(w_{t})\big)+\frac{2}{1-\beta}\left(\alpha_{t}\,\big(F(w_{*})-F(w_{t})\big)+\xi_{t}\right)+\varphi_{t}$
where $\xi_{t}$ and $\varphi_{t}$ are defined as $\mathscr{F}_{t}$-adapted processes
$\xi_{t}$ $=(1-\beta)k_{4}\alpha_{t}\langle\delta m_{t},\,v_{t}\rangle+\alpha_{t}\langle\delta m_{t},\,w_{t}-w_{*}\rangle.$ (4.11)
$\varphi_{t}$ $=\alpha_{t}^{2}\left((a_{1}k_{0}^{2}+k_{1})\,\ell(Z_{t},w_{t})+a_{2}k_{0}^{2}+k_{2}\right).$
By lemma 4.1, we have
$\sum_{t=1}^{\infty}\varphi_{t}\leq(a_{1}k_{0}^{2}+k_{1})K_{2}+(a_{2}k_{0}^{2}+k_{2})<\infty~~~\text{a.s.-}\mathbb{P}.$ (4.12)
Also, we have $\mathbb{E}[\xi_{t}\,|\,\mathscr{F}_{t}]=0$. We now show that we may bound $\sum_{k=1}^{t}\xi_{k}-\mathbb{E}[\xi_{k}\,|\,\mathscr{F}_{k}]$ using [8, lemma 21]. By lemma C.6.3b, lemma 4.1, and (4.9),
$\xi_{t}-\mathbb{E}[\xi_{t}\,|\,\mathscr{F}_{t}]=\xi_{t}$ $\leq\alpha_{t}\|\delta m_{t}\|\left(k_{4}\,\|v_{t}\|+\|w_{t}-w_{*}\|\right)$ (4.13)
$\leq\alpha_{t}\sqrt{6L^{2}\|w_{t}-w_{*}\|^{\gamma}+a_{4}}\left(k_{4}+1\right)\sqrt{\sum_{s=0}^{t-1}\alpha_{s}}$
$\leq\left(k_{4}+1\right)\,\left(6L^{2}\,\sqrt{\alpha_{t}^{1-\gamma}\left(\alpha_{t}\sum_{s=0}^{t-1}\alpha_{s}\right)^{\gamma+1}}+a_{4}\,\sqrt{\alpha_{t}^{2}\sum_{s=0}^{t-1}\alpha_{s}}\right)$
$\leq\left(k_{4}+1\right)\left(6L^{2}\,\sqrt{K_{0}^{1+\gamma}}+a_{4}\sqrt{K_{0}}\right).$
As for the conditional variance, we apply Cauchy-Schwarz, lemma C.6.3d, and lemma 4.1 to get
$\mathbb{E}[\xi_{t}^{2}\,|\,\mathscr{F}_{t}]$ $\leq 2\alpha_{t}^{2}\left(k_{4}^{2}\,\mathbb{E}[\langle\delta m_{t},\,v_{t}\rangle^{2}\,|\,\mathscr{F}_{t}]+\mathbb{E}[\langle\delta m_{t},\,w_{t}-w_{*}\rangle^{2}\,|\,\mathscr{F}_{t}]\right)$
$\leq 2\left(k_{4}^{2}+1\right)\alpha_{t}^{2}\,\mathbb{E}[\|\delta m_{t}\|^{2}\,|\,\mathscr{F}_{t}]\bigg(\|v_{t}\|^{2}+\|w_{t}-w_{*}\|^{2}\bigg)$
$\leq 2\left(k_{4}^{2}+1\right)\alpha_{t}^{2}\bigg(a_{6}\,(F(w_{t})-F_{*})+a_{7}\bigg)\left(\left(\sum_{s=0}^{t-1}\alpha_{s}\right)+\|w_{t}-w_{*}\|^{2}\right).$
Recall, by lemma 4.1, we get $\alpha_{t}\|w_{t}-w_{*}\|^{2}\leq\alpha_{t}\sum_{s=0}^{t-1}\alpha_{s}\leq K_{0}$. Let $k_{11}=2\left(k_{4}^{2}+1\right)$. There exists $c>0$ such that
$\mathbb{E}[\xi_{t}^{2}\,|\,\mathscr{F}_{t}]$ $\leq 2k_{11}\,a_{6}\,K_{0}\,\alpha_{t}\,(F(w_{t})-F_{*})+k_{11}\,K_{0}\,a_{7}+k_{11}\,a_{7}\,\alpha_{t}^{2}\,\|w_{t}-w_{*}\|^{2}$ (4.14)
$\leq c\left(1+\alpha_{t}\,(F(w_{t})-F_{*})+\alpha_{t}^{2}\,\|w_{t}-w_{*}\|^{2}\right).$
Let $b>0$ be the right hand side of (4.13). Applying [8, lemma 21.b], we get for a choice of tolerance level $\delta\in(0,1)$ and all $q>0$, the following holds with probability at least $1-\delta$:
$\sum_{s=1}^{t-1}\xi_{s}$ $\leq\frac{(e^{q}-q-1)}{q}\cdot\frac{c}{b}\left(1+\sum_{s=1}^{t-1}\alpha_{s}\,(F(w_{s})-F_{*})+\sum_{s=1}^{t-1}\alpha_{s}^{2}\,\|w_{s}-w_{*}\|^{2}\right)+\frac{b}{q}\cdot\frac{\log(1/\delta)}{q}.$ (4.15)
Now, we want to ‘remove’ the $\alpha_{t}\,(F(w_{*})-F(w_{t}))$ term in (4.10) so that the upper bound of $\Delta z_{t+1}+k_{0}^{2}\Delta v_{t+1}$ is of the order $O(1+\sum_{k=0}^{t-1}\alpha_{k}^{2}\|w_{k}-w_{*}\|^{2})$. To do so, we choose $q>0$ such that
$\frac{e^{q}-q-1}{q}\cdot\frac{c}{b}<1$ (4.16)
The above inequality is solvable since $(e^{q}-q-1)/q\rightarrow 0$ as $q\rightarrow 0$. Moreover, $q$ is independent of $t$ and $\alpha_{t}$ since $b$ and $c$ are both independent of $t$ and $\alpha_{t}$. Therefore, all constants on the right hand side of (4.15) are independent of $t$ and $\alpha_{t}$. Since $F(w_{t})>F_{*}$, the term $\alpha_{t}(F_{*}-F(w_{t}))$ in (4.10) dominates over that in (4.15). Using the estimate in (4.15) and summing (4.10) from $1$ to t1t-1 yields the desired result. ∎
Remark 4.4.
Observe that the constant $K_{5}$ is independent of the step sizes. This can be seen from (4.16). The constants $b$ and $c$ on the right hand side of (4.13) and (4.14) respectively are independent of the step sizes $\alpha_{t}$.
Proposition 4.5.
With the notation in proposition 4.3, if all assumptions in lemma 4.3 holds and $\alpha_{0}^{2}\leq\frac{1}{4K_{5}}$, then for any $\delta\in(0,1)$ there exists a constant $K_{6}(L,\gamma,\beta)>0$ such that for any $T>0$
$\mathbb{P}\left(\max_{1\leq t\leq T}\left\{\|w_{t}-w_{*}\|^{2},\|z_{t}-w_{*}\|^{2},\|v_{t}\|^{2}\right\}\leq K_{6}\log\left(\frac{T}{\delta}\right)\right)\geq 1-\delta.$
Proof.
It suffices to prove the conclusion for $\|w_{t}-w_{*}\|^{2}$ since $\|z_{t}-w_{*}\|^{2}\leq 2\|w_{t}-w_{*}\|^{2}+2k_{0}^{2}\|v_{t}\|^{2}$ and $v_{t}=(w_{t}-w_{*})-(w_{t-1}-w_{*})$ gives the following inequalities
[…]
As a consequence of proposition 4.3 and $\|w_{t}-w_{*}\|^{2}\leq 2\|z_{t}-w_{*}\|^{2}+2k_{0}^{2}\|v_{t}\|^{2}$, we have
$\mathbb{P}\left(\frac{\|w_{t}-w_{*}\|^{2}}{2}>K_{3}+K_{4}\log\left(\frac{1}{\delta}\right)+K_{5}\sum_{s=1}^{t-1}\alpha_{s}^{2}\,\|w_{s}-w_{*}\|^{2}\right)<\delta~~~,~~~\forall t\geq 1.$ (4.17)
Let $t_{1}=\min\{t\in[1,T]:K_{5}\sum_{s=t}^{T}\alpha_{s}^{2}\leq\frac{1}{4}\}$. Note that $t_{1}$ is well-defined since $\alpha_{T}^{2}\leq\alpha_{0}^{2}\leq\frac{1}{4K_{5}}$. In particular, $t_{1}$ depends only on $L,\gamma,\beta$ since $K_{5}$ depends only on those parameters. Using lemma 4.1, with probability greater than $1-\delta$, for any $t\geq 1$,
$\frac{\|w_{t}-w_{*}\|^{2}}{2}$ $\leq K_{3}+K_{4}\,\log\left(\frac{T}{\delta}\right)+K_{5}\,\sum_{s=1}^{t_{1}}\alpha_{s}^{2}\,\|w_{s}-w_{*}\|^{2}+K_{5}\,\sum_{s=t_{1}}^{T}\alpha_{s}^{2}\,\|w_{s}-w_{*}\|^{2}$
$\leq K_{3}+K_{4}\,\log\left(\frac{T}{\delta}\right)+K_{5}\sum_{s=1}^{t_{1}}\alpha_{s}^{2}\left(\sum_{k=1}^{s}\alpha_{k}\right)+\frac{1}{4}\,\max_{t_{1}\leq t\leq T}\|w_{t}-w_{*}\|^{2}$
$\leq\left(K_{3}+K_{5}\,K_{0}\,\sum_{s=1}^{t_{1}}\alpha_{s}\right)+K_{4}\,\log\left(\frac{T}{\delta}\right)+\frac{1}{4}\,\max_{1\leq t\leq T}\|w_{t}-w_{*}\|^{2}.$
Since the right hand side is independent of $t$, we know $\frac{1}{2}\max_{1\leq t\leq T}\|w_{t}-w_{*}\|^{2}$ is also bounded by the right hand side. Solving the inequality for $\max_{1\leq t\leq T}\|w_{t}-w_{*}\|^{2}$ yields the desired result, where $K_{6}=\max\{4K_{3}+t_{1}K_{0},4K_{4},10k_{0}^{2},10\}$ for all $t>0$. ∎
4.2. Error Rate with High Probability
Corollary 4.6.
With the notation in proposition 4.5, define the event $\pazocal{E}_{t}^{\delta}$ as
$\pazocal{E}_{t}^{\delta}:=\left\{\max\left(\|w_{t}-w_{*}\|^{2},\|z_{t}-w_{*}\|^{2},\|v_{t}\|^{2}\right)\leq K_{6}\,\log\frac{kT}{\delta}\right\}.$
If all assumptions in proposition 4.5 are satisfied, then under the event $\pazocal{E}_{t}^{\delta}$, with the notation in proposition 4.5 and lemma C.6,
  1. 1.
    $\pazocal{E}_{t}^{\delta}\in\mathscr{F}_{t}\subset\mathscr{F}_{T}$ and $\mathbb{P}(\pazocal{E}_{t}^{\delta})\geq 1-\delta$.
  2. 2.
    $F(w_{t})\leq(F_{*}+\frac{L}{1+\gamma}\cdot K_{6})\left(\log\frac{T}{\delta}\right)^{\gamma}$ for all $t\leq T$ and $T>e\delta$.
  3. 3.
    $\|\delta m_{t}\|^{2}\leq(6L^{2}K_{6}+a_{4})\left(\log\frac{T}{\delta}\right)^{\gamma}$ for all $t\leq T$ and $T>e\delta$.
  4. 4.
    $\|\nabla\ell(Z_{t},w_{t})\|^{2}\leq(2L^{2}K_{6}+a_{5})\left(\log\frac{T}{\delta}\right)^{\gamma}$ for all $t\leq T$ and $T>e\delta$.
Proof.
Under the assumptions of corollary 4.6, $T>0$ is fixed and $t\leq T$. Therefore, $\pazocal{E}_{k}$ is in $\mathscr{F}_{t}$. The proof follows directly from proposition 4.5, lemma C.6.2, C.6.3b, and C.6.3c. ∎
Proposition 4.7.
With the notation in lemma 4.1, if all of the assumptions in proposition 4.5 holds, then there exists a constant $K_{7}(L,\gamma,\beta,w_{0},w_{*})>0$ such that for any $T>0$ and $0<\delta<\min\left(1,2T/e\right)$,
$\mathbb{P}\left(\sum_{t=1}^{T}\alpha_{t}\,(F(w_{t})-F_{*})\leq K_{7}\left(\log\frac{2T}{\delta}\right)^{3/2}\right)\geq 1-\delta.$
Proof.
Using the notation in (4.1), rearranging terms in (4.10) and using $F(w_{*})=F_{*}$ gives us
$\frac{2}{1-\beta}\,\alpha_{t}(F(w_{t})-F_{*})\leq 2\,k_{4}\,(\alpha_{t-1}F(w_{t-1})-\alpha_{t}F(w_{t}))-\Delta z_{t+1}-k_{0}^{2}\Delta v_{t+1}+\frac{2\,\xi_{t}}{1-\beta}+\varphi_{t}$ (4.18)
where $\xi_{t}$ and $\varphi_{t}$ are defined in (4.11). By lemma C.6.3b, we have
$\xi_{t}$ $=(1-\beta)k_{4}\,\alpha_{t}\langle\delta m_{t},\,v_{t}\rangle+\,\alpha_{t}\langle\delta m_{t},\,w_{t}-w_{*}\rangle$ (4.19)
$\leq\left(k_{4}+1\right)\,\alpha_{t}\,\|\delta m_{t}\|\,\left(\|v_{t}\|+\|w_{t}-w_{*}\|\right)$
$\leq\left(k_{4}+1\right)\alpha_{t}\,\sqrt{6L^{2}\|w_{t}-w_{*}\|^{2\gamma}+a_{4}}\,(\|v_{t}\|+\|w_{t}-w_{*}\|)$
Let $\pazocal{E}_{t}^{\delta/2}$ be the event in corollary 4.6. For a fixed $\delta\in(0,1)$ and $T\geq t$, define $c=4\sqrt{6L^{2}+a_{4}}\left(k_{4}+1\right)$ and $\xi_{t}^{\prime}=\xi_{t}\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}$. Since $2T/\delta>e$, we have $\|w_{t}-w_{*}\|^{2\gamma}\leq\log(2T/\delta)\leq(\log(2T/\delta))^{2}$. Therefore, by (4.19), we get
$\xi_{t}^{\prime}$ $\leq 2\sqrt{2}\left(k_{4}+1\right)\sqrt{2(6L^{2}+a_{4})\,\alpha_{t}^{2}\cdot\left(\log\frac{2T}{\delta}\right)^{2}}.$
Applying [8, lemma 21.a] to $\xi_{t}^{\prime}$ yields
$\sum_{t=1}^{T}\xi_{t}^{\prime}\leq\left(c\log\frac{2T}{\delta}\right)\left(2\log\frac{1}{\delta}\sum_{t=1}^{T}\alpha_{t}^{2}\right)^{\frac{1}{2}}\leq c\sqrt{2}\left(\log\frac{2T}{\delta}\right)^{\frac{3}{2}}.$ (4.20)
Let $K^{\prime}>0$ be the constant on the right hand side of (4.12). Applying proposition 4.5 and substituting (4.20) to (4.18) yields that the following occurs with probability at least $1-\delta$:
$\sum_{t=1}^{T}\alpha_{t}(F(w_{t})-F_{*})\leq\frac{1-\beta}{2}\left(\alpha_{0}F(w_{0})+\|z_{1}-w_{*}\|^{2}+\|v_{1}\|^{2}+\frac{2}{1-\beta}\cdot c\sqrt{2}\left(\log\frac{2T}{\delta}\right)^{\frac{3}{2}}+K^{\prime}\right).$
Since $z_{1}=w_{0}$, $v_{1}=0$, and $2T/\delta>e$, we take $K_{7}=4\max\left(\alpha_{0}\,F(w_{0}),\|w_{0}-w_{*}\|^{2},c\sqrt{2},K^{\prime}\right)$. ∎
Proposition 4.8.
With the notation as in lemma 4.1, if all assumptions in proposition 4.5 holds, then there exists a constant $K_{8}(L,\gamma,\beta)>0$ such that for any $\delta\in(0,1)$, $T>1$, and $\tau\in[1,T]$,
$\mathbb{P}\left[\sum_{t=\tau+1}^{T}\alpha_{t}\,\|v_{t}\|^{2}\leq K_{8}\cdot\max\left(\alpha_{\tau},\sqrt{\sum_{t=\tau}^{T}\alpha_{t}^{4}},\sum_{t=\tau}^{T}\alpha_{t}^{3}\right)\left(\log\frac{2T}{\delta}\right)^{2}\right]\geq 1-\delta.$
Proof.
Consider the recursion $\alpha_{t+1}\,\|v_{t+1}\|^{2}\leq\beta^{2}\alpha_{t}\,\|v_{t}\|^{2}-2\alpha_{t}^{2}\beta\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle+\alpha_{t}^{3}\|\nabla\ell(Z_{t},w_{t})\|^{2}$ obtained from the recursion of $\|v_{t}\|^{2}$. Since $v_{1}=0$, $F$ is convex and $\alpha_{t}$ is decreasing, summing both sides from $t=\tau$ to $t=T$ yields
$(1-\beta^{2})\sum_{t=\tau+1}^{T}\alpha_{t}\,\|v_{t}\|^{2}$ $\leq\beta^{2}\alpha_{\tau}\|v_{\tau}\|^{2}-\alpha_{T+1}\|v_{T+1}\|^{2}+\sum_{t=\tau}^{T}-2\beta\alpha_{t}^{2}\bigg(\langle\nabla F(w_{t}),\,v_{t}\rangle-\langle\delta m_{t},\,v_{t}\rangle\bigg)$
$+\sum_{t=\tau}^{T}\alpha_{t}^{3}\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq\beta^{2}\alpha_{\tau}\|v_{\tau}\|^{2}+2\beta(\alpha_{\tau}F(w_{\tau})-\alpha_{T}F(w_{T}))$
$+\sum_{t=\tau}^{T}\bigg\{2\beta\alpha_{t}^{2}\langle\delta m_{t},\,v_{t}\rangle+\alpha_{t}^{3}\|\nabla\ell(Z_{t},w_{t})\|^{2}\bigg\}.$
By corollary 4.6, there exists an event $\pazocal{E}_{t}^{\delta/2}$ such that
$\sum_{t=\tau+1}^{T}\alpha_{t}\,\|v_{t}\|^{2}\leq c_{0}\cdot\left(\log\frac{2T}{\delta}\right)^{2}\left(\alpha_{\tau}+\sum_{t=\tau}^{T}\left\{\alpha_{t}^{2}\langle\delta m_{t},\,v_{t}\rangle+\alpha_{t}^{3}\right\}\right).$ (4.21)
where
$c_{0}=\frac{1}{1-\beta^{2}}\cdot\max\left(\beta^{2},2\beta\left(F_{*}+\frac{LK_{6}}{1+\gamma}\right),2L^{2}K_{6}+a_{5}\right).$
It remains to estimate the martingale part in (4.21). Applying Cauchy-Schwarz and lemma C.6.3b gives
$\alpha_{t}^{2}\cdot|\langle\delta m_{t},\,v_{t}\rangle|\cdot\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}\leq\alpha_{t}^{2}\cdot\sqrt{6L^{2}+a_{4}}\cdot\left(\log\frac{2T}{\delta}\right)^{\gamma}$
Using Azuma-Hoeffding [8, lemma 21.a], there exists an event $\pazocal{E}\in\mathscr{F}_{T}$ such that $\mathbb{P}(\pazocal{E})\geq 1-\delta$ and $\cap_{t=\tau}^{T}\pazocal{E}_{t}^{\delta/2}\subset\pazocal{E}$ for which the following occurs
$\sum_{t=\tau+1}^{T}\alpha_{t}\|v_{t}\|^{2}\leq c_{0}\cdot\left(\log\frac{2T}{\delta}\right)^{2}\left(\alpha_{\tau}+\sqrt{2\sum_{t=\tau}^{T}\alpha_{t}^{4}}+\sum_{t=\tau}^{T}\alpha_{t}^{3}\right).$
This completes the proof. ∎
Lemma 4.9.
With the notation as in lemma 4.1, if all assumptions in proposition 4.5 holds, then there exists a constant $K_{9}(L,\gamma,\beta)>0$ such that for any $\delta\in(0,1)$, $T>1$, and $\tau\in[1,T]$,
$\mathbb{P}\left[\sum_{t=\tau}^{T}\alpha_{t}\langle\delta m_{t},\,v_{t}\rangle\leq K_{9}\cdot\max\left(\alpha_{\tau-1},\sqrt{\sum_{t=\tau-1}^{T}\alpha_{t}^{4}},\sum_{t=\tau-1}^{T}\alpha_{t}^{3}\right)\left(\log\frac{3T}{\delta}\right)^{2}\right]\geq 1-\delta.$
Proof.
Let $\pazocal{E}_{t}^{\delta/3}$ be the event defined in corollary 4.6. Using Bernstein’s inequality [8, lemma 21.b], we can bound the martingale $\sum_{t=\tau}^{T}\alpha_{t}\,\langle\delta m_{t},\,v_{t}\rangle$ from above. By corollary 4.6, we have the estimates
$\alpha_{t}\cdot|\langle\delta m_{t},\,v_{t}\rangle|\cdot\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}\leq\sqrt{6L^{2}K_{6}+a_{4}}\cdot\alpha_{\tau}\cdot\left(\log\frac{3T}{\delta}\right)^{\gamma}$
and
$\alpha_{t}^{2}\cdot\mathbb{E}[|\langle\delta m_{t},\,v_{t}\rangle|^{2}\cdot\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}\,|\,\mathscr{F}_{t}]\leq(6L^{2}K_{6}+a_{4})\cdot\alpha_{\tau}\cdot\alpha_{t}\cdot\|v_{t}\|^{2}\cdot\left(\log\frac{3T}{\delta}\right)^{\gamma}.$
Let $b=\sqrt{6L^{2}K_{6}+a_{4}}$. Choosing $q>0$ such that $(e^{q}-q-1)/q=1$, by Bernstein’s inequality [8, lemma 21.b], the following occurs with probability at least $1-\delta/3$:
$\sum_{t=\tau}^{T}\alpha_{t}\,\langle\delta m_{t},\,v_{t}\rangle\leq b\left(\sum_{t=\tau}^{T}\alpha_{t}\,\|v_{t}\|^{2}\right)+\frac{b}{q}\cdot\alpha_{\tau}\left(\log\frac{3}{\delta}\right).$ (4.22)
Let $\pazocal{E}^{\prime}\in\mathscr{F}_{T}$ denote the event of (4.22) and $\pazocal{E}_{T}^{\prime}\in\mathscr{F}_{T}$ be the event in the conclusion of proposition 4.8 replacing $\delta$ by $2\delta/3$ instead. Note that $\mathbb{P}(\pazocal{E}^{\prime}\cap\pazocal{E}_{T}^{\prime})\geq 1-\delta$ and in the event $\pazocal{E}^{\prime}\cap\pazocal{E}_{T}^{\prime}\in\mathscr{F}_{T}$,
$\sum_{t=\tau}^{T}\alpha_{t}\,\langle\delta m_{t},\,v_{t}\rangle\leq b\cdot\left(K_{8}+\frac{1}{q}\right)\cdot\max\left(\alpha_{\tau-1},\sqrt{\sum_{t=\tau-1}^{T}\alpha_{t}^{4}},\sum_{t=\tau-1}^{T}\alpha_{t}^{3}\right)\cdot\left(\log\frac{3T}{\delta}\right)^{2}.$
This completes the proof. ∎
Lemma 4.10.
With the notation as in lemma 4.1, if all assumptions in proposition 4.5 holds, then there exists a constant $K_{10}(L,\gamma,\beta)>0$ such that for any $\delta\in(0,1)$, $T>1$, and $\tau\in[0,T]$,
$\mathbb{P}\left[\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,-\alpha_{t}\nabla\ell(Z_{t},w_{t})\rangle\leq K_{10}\cdot\alpha_{\tau}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}\right]\geq 1-\delta.$
Proof.
We may decompose $\alpha_{t}\langle\nabla F(w_{t}),\,\nabla\ell(Z_{t},w_{t})\rangle=-\alpha_{t}\|\nabla F(w_{t})\|^{2}+\psi_{t}$ where we define
$\psi_{t}^{\prime}=\psi_{t}\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}~,~\psi_{t}=\alpha_{t}\langle\nabla F(w_{t}),\,\delta m_{t}\rangle.$ (4.23)
By Cauchy-Schwarz, $\gamma$-Hölder property of $\nabla F$, lemma C.6.3b, for all $t\in[\tau,T]$,
$|\psi_{t}^{\prime}|$ $\leq\mathbbm{1}_{\pazocal{E}}\cdot\alpha_{t}\cdot\|\nabla F(w_{t})\|\cdot\|\delta m_{t}\|$ (4.24)
$\leq L\mathbbm{1}_{\pazocal{E}}\cdot\alpha_{t}\sqrt{6L^{2}\|w_{t}-w_{*}\|^{4\gamma}+a_{4}\|w_{t}-w_{*}\|^{2\gamma}}$
$\leq L\sqrt{K_{6}^{\gamma}(6L^{2}K_{6}^{\gamma}+a_{4})}\cdot\alpha_{\tau}\cdot\left(\log\frac{2T}{\delta}\right)^{2\gamma}.$
We apply Cauchy-Schwarz and lemma C.6.3b to get
$\mathbb{E}[(\psi_{t}^{\prime})^{2}\,|\,\mathscr{F}_{t}]$ $\leq\alpha_{t}^{2}\cdot\|\nabla F(w_{t})\|^{2}\cdot\mathbb{E}[\|\delta m_{t}\|^{2}\cdot\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}\,|\,\mathscr{F}_{t}]$ (4.25)
$\leq\alpha_{\tau}\cdot\alpha_{t}\,\|\nabla F(w_{t})\|^{2}\cdot(6L^{2}\|w_{t}-w_{*}\|^{2\gamma}\mathbbm{1}_{\pazocal{E}_{t}^{\delta/2}}+a_{4})$
$\leq\alpha_{\tau}\cdot\alpha_{t}\,\|\nabla F(w_{t})\|^{2}\cdot(6L^{2}+a_{4})\cdot\left(\log\frac{2T}{\delta}\right)^{\gamma}.$
For shorthand notation, denote by $b=L\sqrt{K_{6}^{\gamma}(6L^{2}K_{6}+a_{4})}$ and $c=6L^{2}+a_{4}$ as the constants appearing on the right hand side of (4.24) and (4.25). By Bernstein’s inequality [8, lemma 21.b], for any $q>0$ and $\delta\in(0,1)$, there exists an event $\pazocal{E}^{\prime}\in\mathscr{F}_{T}$ with $\mathbb{P}(\pazocal{E}^{\prime})\geq 1-\frac{\delta}{2}$ where $\pazocal{E}^{\prime}$ is the event
$\sum_{t=\tau}^{T}\psi_{t}^{\prime}\leq\frac{e^{q}-q-1}{q}\cdot\frac{c}{b}\left(\sum_{t=\tau}^{T}\alpha_{t}\,\|\nabla F(w_{t})\|^{2}\right)+\alpha_{\tau}\cdot\frac{b}{q}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}.$ (4.26)
Since $(e^{q}-q-1)/q\rightarrow 0$ as $q\rightarrow 0$, may choose $q>0$ such that
$\frac{e^{q}-q-1}{q}\cdot\frac{c}{b}<1.$
Let $\pazocal{E}\in\mathscr{F}_{T}$ be the event in the conclusion of proposition 4.5, except with $\delta$ replaced by $\delta/2$. Under the event $\pazocal{E}\cap\pazocal{E}^{\prime}$, we know $\psi_{t}^{\prime}=\psi_{t}$ and the following occurs with probability $\mathbb{P}(\pazocal{E}\cap\pazocal{E}^{\prime})\geq 1-\delta$
$\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,-\alpha_{t}$ $\nabla\ell(Z_{t},w_{t})\rangle=\sum_{t=\tau}^{T}\left(-\alpha_{t}\,\|\nabla F(w_{t})\|^{2}+\psi_{t}^{\prime}\right)$
$\leq-\left(\left(1-\frac{e^{q}-q-1}{q}\cdot\frac{c}{b}\right)\sum_{t=\tau}^{T}\alpha_{t}\,\|\nabla F(w_{t})\|^{2}\right)+\frac{b}{q}\cdot\alpha_{\tau}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}$
$\leq\frac{b}{q}\cdot\alpha_{\tau}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}.$
The proof is complete. ∎
Proposition 4.11.
With the notation in lemma 4.1, if all assumptions of proposition 4.5 holds, $\beta\in[0,1)$, and there exists $c_{1},c_{2}\in\mathbb{R}_{+}$ such that $c_{1}\,t^{-p}\leq\alpha_{t}\leq c_{2}\,t^{-p}$ where $p\in(\frac{1}{2},1)$, then there exists a constant $K_{11}(L,\beta,\gamma,w_{0},\|\alpha_{t}\|_{\ell^{2}(\mathbb{N})})>0$ such that for any $\delta\in(0,1)$, $T>1$, and $\tau\in[1,T]$,
$\mathbb{P}\left[\sum_{t=\tau}^{T}\|v_{t+1}\|^{2}\leq K_{11}\,\left(\log\frac{3T}{\delta}\right)^{2}\max\left(\tau^{-2p+1},\,\beta^{2\tau}\right)\right]\geq 1-\delta.$
Proof.
By lemma 4.9 and proposition 4.5, there exists an event $\pazocal{E}_{T}^{\prime}\in\mathscr{F}_{T}$ such that with probability $1-\delta$ both the conclusion in lemma 4.9 and proposition 4.5 holds. We have the recursion $\|v_{t+1}\|^{2}=\beta^{2}\|v_{t}\|^{2}-2\beta\alpha_{t}\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle+\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}$.
Case $\beta=0$. If $\beta=0$, then $\|v_{t+1}\|^{2}=\alpha_{t}^{2}\|\nabla\ell(Z_{t},w_{t})\|^{2}$. Under the event $\pazocal{E}_{T}^{\prime}$, then by corollary 4.6, we have
$\sum_{t=\tau}^{T}\|v_{t+1}\|^{2}=\sum_{t=\tau}^{T}\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}\lesssim\left(\log\frac{T}{\delta}\right)^{\gamma}\sum_{t=\tau}^{T}\alpha_{t}^{2}\leq\left(\log\frac{T}{\delta}\right)^{\gamma}\tau^{-2p+1}.$
Case $\beta\in(0,1)$. The recursion of $\|v_{t}\|^{2}$ in the previous paragraph can be rewritten into an integral equation by proposition B.1:
$\begin{cases}\|v_{t+1}\|^{2}=\sum_{s=0}^{t-1}\beta^{2s}X_{t-s}.\\ X_{t}=-2\beta\alpha_{t}\langle\nabla\ell(Z_{t},w_{t}),\,v_{t}\rangle+\alpha_{t}^{2}\|\nabla\ell(Z_{t},w_{t})\|^{2}.\end{cases}$
It follows that
$\sum_{t=\tau}^{T}\|v_{t+1}\|^{2}$ $=\sum_{t=\tau}^{T}\sum_{s=1}^{t}\beta^{2(s-1)}X_{t-s+1}$ (4.27)
$=\left(\sum_{s=1}^{\tau}\sum_{t=\tau}^{T}+\sum_{s=\tau}^{T}\sum_{t=s}^{T}\right)\beta^{2(s-1)}X_{t-s+1}$
$=\sum_{s=1}^{\tau}\beta^{2(s-1)}\left(\sum_{t=\tau}^{T}X_{t-s+1}\right)+\beta^{2(\tau-1)}\sum_{s=0}^{T-\tau}\beta^{2s}\left(\sum_{t=1}^{T-\tau-s+1}X_{t}\right).$
We estimate the parentheses in the first term. By lemma C.6.3c and the convexity of $F$, we get
$X_{t}$ $=-2\beta\,\alpha_{t}\,\langle\nabla F(w_{t}),\,v_{t}\rangle+2\,\alpha_{t}\,\beta\langle\delta m_{t},\,v_{t}\rangle+\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}$ (4.28)
$\leq 2\beta\alpha_{t}\left(F(w_{t-1})-F(w_{t})+\langle\delta m_{t},\,v_{t}\rangle\right)+\alpha_{t}^{2}\,(2L^{2}\|w_{t}-w_{*}\|^{2\gamma}+a_{5})$
$\leq 2\beta(\alpha_{t-1}F(w_{t-1})-\alpha_{t}F(w_{t}))+2\beta\alpha_{t}\langle\delta m_{t},\,v_{t}\rangle+\alpha_{t}^{2}\,(2L^{2}\|w_{t}-w_{*}\|^{2\gamma}+a_{5}).$
Using corollary 4.6 and $\alpha_{t}=\Theta(t^{-p})$ with $p\in(\frac{1}{2},1)$, the following holds in $\pazocal{E}_{T}^{\prime}$
$\sum_{t=\tau}^{T}X_{t-s+1}$ $\lesssim\alpha_{\tau-s}F(w_{\tau-s})+\sum_{t=\tau-s+1}^{T-s+1}\alpha_{t}\langle\delta m_{t},\,v_{t}\rangle+\left(\log\frac{2T}{\delta}\right)^{\gamma}\sum_{s\leq t=\tau}^{T}\alpha_{t-s+1}^{2}$ (4.29)
$\lesssim\left(\log\frac{3T}{\delta}\right)^{2}\left(2\alpha_{\tau-s}+(\tau-s+1)^{-2p+1}\right)$
$\lesssim\left(\log\frac{3T}{\delta}\right)^{2}(\tau-s)^{-2p+1}.$
Therefore, under the event $\pazocal{E}_{T}^{\prime}$,
$\sum_{s=1}^{\tau}\sum_{t=\tau}^{T}\beta^{2(s-1)}X_{t-s+1}\lesssim\left(\log\frac{3T}{\delta}\right)^{2}\tau^{-2p+1}\sum_{s=1}^{\tau}\beta^{2(s-1)}\left(\frac{\tau}{\tau-s}\right)^{2p-1}.$
Now, the ratio on the right hand side converges to 1 as $\tau\uparrow\infty$. As a consequence, the series $\sum_{s=1}^{\tau}\beta^{2(s-1)}(\tau/(\tau-s))^{2p-1}$ converges as $\tau\uparrow\infty$ by the dominated convergence theorem. Therefore, such a series in (4.2) is bounded by a constant independent of $\tau$.
Now, examine the second term on the right hand side of (4.27). By lemma C.6.3a, we have $\alpha_{t}^{2}\,\|\nabla\ell(Z_{t},w_{t})\|^{2}\leq\alpha_{t}^{2}\,(a_{1}\ell(Z_{t},w_{t})+a_{2})$. By lemma 4.1, we get $\sum\alpha_{t}^{2}\,\ell(Z_{t},w_{t})<K_{2}$. Therefore, using the expression for $X_{t}$ in (4.28), under the event $\pazocal{E}_{T}^{\prime}$,
$\sum_{t=1}^{T-\tau-s+1}X_{t}\lesssim F(w_{0})+\left(\log\frac{3T}{\delta}\right)^{2}\left(\max\left(1,\sqrt{\sum_{t=1}^{\infty}\alpha_{t}^{4}},\sum_{t=1}^{\infty}\alpha_{t}^{3}\right)+\left(a_{1}\,K_{2}+a_{2}\sum_{t=0}^{\infty}\alpha_{t}^{2}\right)\right).$
Therefore, the claim is proved. ∎
Proposition 4.12.
With the notation in lemma 4.1, if all the assumptions in proposition 4.11 holds and $\gamma=1$, then there exists a constant $K_{12}(L,\beta,\gamma,w_{0},\|\alpha_{t}\|_{\ell^{2}(\mathbb{N})})>0$ such that for any $\delta\in(0,1)$, $T>1$, and $\tau\in(1,T]$,
$\mathbb{P}\left[\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,v_{t}\rangle\leq K_{12}\,\left(\log\frac{3T}{\delta}\right)^{2}\max\left(\tau^{-2p+1},\,\beta^{2\tau}\right)\right]\geq 1-\delta.$
Proof.
Fix $T>1$ and $\tau>1$. Let $\Lambda_{t}=\langle\nabla F(w_{t}),\,-\alpha_{t}\nabla\ell(Z_{t},w_{t})\rangle$. By lemma 4.10 and proposition 4.11, there exists an event $\pazocal{E}\in\mathscr{F}$ such that with probability $\mathbb{P}(\pazocal{E})\geq 1-\delta$ the following occurs:
$\sum_{t=\tau}^{T}\|v_{t+1}\|^{2}\leq K_{11}\cdot\max(\tau^{-2p+1},\beta^{2\tau})\left(\log\frac{3T}{\delta}\right)^{2}~~~,~~~\forall~\tau\in[1,T].$ (4.30)
$\sum_{t=\tau}^{T}\Lambda_{t}\leq K_{10}\cdot\alpha_{\tau}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}~~~,~~~\forall~\tau\in[0,T].$ (4.31)
Since $\gamma=1$, i.e. $\nabla F$ is Lipschitz, $\langle\nabla F(w_{t}),\,v_{t}\rangle$ satisfies the following inequality for any $t>1$,
$\langle\nabla F(w_{t}),\,v_{t}\rangle$ $=\langle\nabla F(w_{t})-\nabla F(w_{t-1})+\nabla F(w_{t-1}),\,v_{t}\rangle$
$=\langle\nabla F(w_{t-1}),\,\beta v_{t-1}-\alpha_{t-1}\nabla\ell(Z_{t-1},w_{t-1})\rangle+\langle\nabla F(w_{t})-\nabla F(w_{t-1}),\,v_{t}\rangle$
$=\beta\langle\nabla F(w_{t-1}),\,v_{t-1}\rangle+\Lambda_{t-1}+\langle\nabla F(w_{t})-\nabla F(w_{t-1}),\,v_{t}\rangle$
$\leq\beta\,\langle\nabla F(w_{t-1}),\,v_{t-1}\rangle+\Lambda_{t-1}+\|v_{t}\|^{2}.$
Also, since $v_{1}=0$, we have $\langle\nabla F(w_{1}),\,v_{1}\rangle=0$. Using proposition B.1 and switching the order of integration, under the event $\pazocal{E}$, the following occurs with probability $\mathbb{P}(\pazocal{E})\geq 1-\delta$ for all $\tau>1$:
$\sum_{t=\tau}^{T}\langle\nabla F(w_{t})$ $,\,v_{t}\rangle\leq\sum_{s=0}^{\tau-2}\beta^{s}\left(\sum_{t=\tau}^{T}\Lambda_{t-s-1}+\|v_{t-s}\|^{2}\right)+\sum_{s=\tau-2}^{T-2}\beta^{s}\left(\sum_{t=s+2}^{T}\Lambda_{t-s-1}+\|v_{t-s}\|^{2}\right)$
$=\sum_{s=0}^{\tau-2}\beta^{s}\left(\sum_{t=\tau}^{T}\Lambda_{t-s-1}+\|v_{t-s}\|^{2}\right)+\beta^{\tau}\sum_{s=0}^{T-\tau}\beta^{s-2}\left(\sum_{t=\tau}^{T}\Lambda_{t-\tau+1}+\|v_{t-\tau+2}\|^{2}\right)$
$\lesssim\left(\log\frac{3T}{\delta}\right)^{2}\left(\sum_{s=0}^{\tau-2}\beta^{s}\cdot\alpha_{\tau-s-1}+\beta^{\tau}\sum_{s=0}^{\infty}\beta^{s-2}\right)$
$\lesssim\max(\tau^{-2p+1},\beta^{\tau})\cdot\left(\log\frac{3T}{\delta}\right)^{2}\left(\sum_{s=0}^{\tau-2}\beta^{s}\left(\frac{\tau}{\tau-s-1}\right)^{p}+\frac{1}{\beta^{2}(1-\beta)}\right).$
The summation on the right hand side remains bounded as $\tau\uparrow\infty$ by the dominated convergence theorem so that the summation term is bounded by a constant $c>0$ independent of $\tau$. The coefficient $K_{12}$ must be of the order $K_{12}=O(c\cdot\max(K_{10},K_{11}))$. ∎
Theorem 4.13.
If all assumptions in proposition 4.5 and proposition 4.12 both holds, then there exists a constant $K_{13}(L,\beta,\gamma,w_{0},w_{*},\|\alpha_{t}\|_{\ell^{2}(\mathbb{N})})>0$ such that for all $T>1$, $T_{0}\in(1,T]$, $\delta\in(0,1)$
$\mathbb{P}\left(F(w_{T+1})-F_{*}\leq K_{13}\left(\log\frac{3T}{\delta}\right)^{2}\,\max\left\{\frac{1}{\sum_{t=T_{0}}^{T}\alpha_{t}}\,,\,T_{0}^{-2p+1}\,,\,\beta^{T_{0}}\right\}\right)\geq 1-\delta.$
Proof.
Let $\Delta_{t}F=F(w_{t})-F_{*}$. By proposition 4.7, proposition 4.11, lemma 4.10, and proposition 4.12, there exists an event $\pazocal{E}\in\mathscr{F}_{T}$ with $\mathbb{P}(\pazocal{E})\geq 1-\delta$ such that all of the following holds
$\sum_{t=1}^{T}\alpha_{t}\,\Delta_{t}F\leq K_{7}\left(\log\frac{2T}{\delta}\right)^{\frac{3}{2}}.$ (4.32)
$\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,-\alpha_{t}\nabla\ell(Z_{t},w_{t})\rangle\leq K_{10}\cdot\alpha_{\tau}\left(\log\frac{2T}{\delta}\right)^{1+\gamma}.$ (4.33)
$\sum_{t=\tau}^{T}\left\{\|v_{t+1}\|^{2}+\langle\nabla F(w_{t}),\,v_{t}\rangle\right\}\leq(K_{11}+K_{12})\left(\log\frac{3T}{\delta}\right)^{2}\max\left(\tau^{-2p+1},\beta^{\tau}\right).$ (4.34)
If we fix $T>1$ and $T_{0}\in(1,T]$, then under the event $\pazocal{E}$, there exists $\tau\in[T_{0},T]$ such that
$\Delta_{\tau}F\leq\sum_{t=T_{0}}^{T}\frac{\alpha_{t}\,\Delta_{t}F}{\sum_{t=T_{0}}^{T}\alpha_{t}}\leq\frac{K_{7}\left(\log\frac{2T}{\delta}\right)^{\frac{3}{2}}}{\sum_{t=T_{0}}^{T}\alpha_{t}}$
where the middle term is the weighted average of $\Delta_{t}F$ from $\tilde{t}$ to $T$. Using $(\gamma,L)$-smoothness of $F$ along with estimates (4.32), (4.33), and (4.34), on the event $\pazocal{E}_{1}$, we have
$\Delta_{T+1}F$ $\leq\Delta_{\tau}F+\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,w_{t+1}-w_{t}\rangle+\frac{L}{1+\gamma}\sum_{t=\tau}^{T}\|v_{t+1}\|^{1+\gamma}$ (4.35)
$=\Delta_{\tau}F-\sum_{t=\tau}^{T}\bigg\{\alpha_{t}\,\|\nabla F(w_{t})\|^{2}-\psi_{t}\bigg\}+\beta\sum_{t=\tau}^{T}\langle\nabla F(w_{t}),\,v_{t}\rangle+\frac{L}{2}\sum_{t=\tau}^{T}\|v_{t+1}\|^{2}.$
$\lesssim K\left(\log\frac{3T}{\delta}\right)^{2}\max\left\{\frac{1}{\sum_{t=T_{0}}^{T}\alpha_{t}},\alpha_{\tau},\max(\tau^{-2p+1},\beta^{\tau})\right\}.$
where
$K=\max\left(K_{7},K_{10},\frac{L+2\beta}{2}(K_{11}+K_{12})\right).$
The proof is complete. ∎
Remark 4.14.
Theorem 2.10 follows from setting $T_{0}=\lfloor T/2\rfloor$. Observe that $\sum_{t=\lfloor T/2\rfloor}^{T}\alpha_{t}\geq\alpha_{T}\,\lfloor\frac{T}{2}\rfloor$. Therefore, setting $T_{0}=\lfloor T/2\rfloor$ in the conclusion of theorem 4.13 yields
$\max\left(T^{-2p+1},T^{p-1},\beta^{T}\right)=O(T^{\max(p-1,-2p+1)})~~~,~~p\in\left(\frac{1}{2},1\right).$
Therefore, theorem 2.10 follows.
Appendix A Proof of Proposition 2.4 and Lemma 2.5
A.1. Proof of Proposition 2.4
Proof.
Since $X_{t}$ is non-negative, we may drop $X_{t}$ from the assumed inequality on $Y_{t}$ and obtain $Y_{t}-Y_{t-1}\leq a_{t-1}Y_{t-1}-X_{t-1}+Z_{t-1}\leq a_{t-1}Y_{t-1}+Z_{t-1}$. We may sum up both sides to obtain the recursive relation
$Y_{t}\leq Y_{0}+\sum_{s=1}^{t}a_{s-1}Y_{s-1}+\sum_{s=1}^{t}Z_{s-1}.$ (A.1)
Since $Z_{t}\in\ell^{1}(\mathbb{N})$, the above is a Gronwall-type inequality. By proposition B.2, we get
$Y_{t}\lesssim(Y_{0}+\|Z\|_{\ell^{1}})\prod_{s=1}^{\infty}(1+a_{s-1}).$
Since $a_{t}\in\ell^{1}$, the right hand side converges. Therefore, the first and second claim are proved. To prove the second claim, repeating the same argument to produce (A.1), but without dropping $X_{t}$, we get
$\sum_{s=1}^{t}X_{s-1}\leq Y_{t}+\sum_{s=1}^{t}X_{s-1}\leq Y_{0}+\sum_{s=1}^{t}a_{s-1}Y_{s-1}+\sum_{s=1}^{t}Z_{s-1}.$
By the first claim, $Y_{t}$ is uniformly bounded. Therefore, the right hand side is finite as $t\rightarrow\infty$. The proof is complete. ∎
A.2. Proof of Lemma 2.5
Proof.
We follow [9, lemma 19]. Let $r\in(0,1)$. By the mean value theorem, there exists $\delta\in(0,1)$ such that $(t+1)^{r}-t^{r}=r(t+\delta)^{r-1}\leq rt^{r-1}$. For shorthand notation, let $r_{\gamma}=\frac{\gamma}{1+\gamma}$ and $a_{1},a_{2}\in\mathbb{R}_{+}$ be such that $a_{1}t^{-p}\leq\alpha_{t}\leq a_{2}t^{-p}$ for all $t>0$. For a fixed $s>0$ and any $t\geq s$, we get
$\mathbb{E}[(t+1)^{r}Y_{t+1}\,|\,\mathscr{F}_{s}]$ $\leq(1+c_{1}\,\alpha_{t}^{1+\gamma})\left(1+r\cdot t^{-1}\right)\,\mathbb{E}[t^{r}Y_{t}\,|\,\mathscr{F}_{s}]$ (A.2)
$+\alpha_{t}^{1+\gamma}\left(t^{r}+r\cdot t^{r-1}\right)\left(\mathbb{E}[Z_{t}\,|\,\mathscr{F}_{s}]+c_{2}\right)$
$\leq(1+c_{1}\,\alpha_{t}^{1+\gamma})\,\mathbb{E}[t^{r}Y_{t}\,|\,\mathscr{F}_{s}]+(1+c_{1}\,\alpha_{t}^{1+\gamma})\,r\,\mathbb{E}[t^{r-1}Y_{t}\,|\,\mathscr{F}_{s}]$
$+a_{2}\left(t^{-p(1+\gamma)+r}+t^{-1-p(1+\gamma)+r}\right)\left(\mathbb{E}[Z_{t}\,|\,\mathscr{F}_{s}]+c_{2}\right).$
We require that r1<pr-1<-p and $p(1+\gamma)+r<\min(-p,-1)$ so that the second and third term on the right hand side are both summable. Since $p\in(\frac{1}{1+\gamma},1)$, such condition holds if and only if $r<\min(1-p,p(1+\gamma)-1)$. Now, applying proposition 2.4 yields $\limsup_{t\rightarrow\infty}\,\mathbb{E}[t^{r}\cdot Y_{t}\,|\,\mathscr{F}_{s}]<\infty$. Note that under the same condition on $r$, we may apply expectation on both sides of (A.2) and then apply proposition 2.4 to conclude $\sup_{t\geq 0}\,\mathbb{E}[t^{r}\cdot Y_{t}]<\infty$.
Rearranging (A.2) and summing up from $t=s$ to t=T1t=T-1 yields two inequalities:
$\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}$ $\lesssim\mathbb{E}\left[\sum_{t=s}^{T-1}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=s}^{T-1}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right].$ (A.3)
$\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}$ $\leq\mathbb{E}\left[\sum_{t=1}^{\infty}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\bigg|\mathscr{F}_{s}\right]+\mathbb{E}\left[\sum_{t=1}^{\infty}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{s}\right].$ (A.4)
Taking absolute value on both sides of (A.3), the limit $T\uparrow\infty$ (with Lebesgue’s monotone convergence theorem since summands are non-negative), and then $s\uparrow\infty$ yields that the right hand side is zero. Therefore, we may conclude
$\lim_{s\uparrow\infty}\left(\lim_{T\uparrow\infty}\bigg|\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}\bigg|\right)=0~~~~\text{a.s.-}\mathbb{P}.$
This implies that $\lim_{s\uparrow\infty}s^{r}\,Y_{s}$ is $\mathscr{F}_{\infty}$-measurable. Furthermore, since $\sum_{t=1}^{\infty}\alpha_{t}\,Y_{t}<\infty$ almost surely, by [7, theorem 3.3.1], we have $\min_{s\leq t}Y_{s}=o(t^{p-1})$.
We need to conclude $\liminf_{s\uparrow\infty}s^{r}Y_{s}=0$ a.s.-$\mathbb{P}$. This holds if we require $Y_{t}>0$ for all $t\geq 0$ (counterexample can be constructed if otherwise) assuming $t^{r}\min_{s\leq t}Y_{s}\downarrow 0$ as $t\uparrow\infty$. If $Y_{t}>0$ for all $t\geq 0$, then one can take a subsequence $Y_{t_{n}}\downarrow 0$ decreasing to zero. For instance, one can take $\{t_{n}:n\in\mathbb{N}\}$ such that $Y_{t_{n}}=\min_{t\leq t_{n}}Y_{t}$. Therefore, we have $t_{n}^{r}Y_{t_{n}}=t_{n}^{r}\min_{t\leq t_{n}}Y_{t}$ where the RHS converges to zero. This implies that there exists a subsequence of $t^{r}Y_{t}$ converging to zero, i.e. $\liminf_{s\uparrow\infty}s^{r}Y_{s}=0$ a.s.-$\mathbb{P}$. Therefore, taking $\liminf_{s\uparrow\infty}$ on both sides of (A.4) yields almost surely
$\lim_{T\uparrow\infty}T^{r}\cdot Y_{T}$ $=\lim_{T\uparrow\infty}\liminf_{s\uparrow\infty}\bigg(\mathbb{E}[T^{r}\cdot Y_{T}\,|\,\mathscr{F}_{s}]-s^{r}\cdot Y_{s}\bigg)$
$\leq\left(\sum_{t=1}^{\infty}(\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}\right)+\mathbb{E}\left[\sum_{t=1}^{\infty}t^{-p(1+\gamma)+r}Z_{t}\bigg|\mathscr{F}_{\infty}\right]$
$<\infty.$
Since $\mathbb{E}[\lim_{t\uparrow\infty}t^{r}Y_{t}\,|\,\mathscr{F}_{s}]$ converges as $s\uparrow\infty$, we know
[…]
If we take $0<s<T$ and $s\uparrow\infty$ so that $T\uparrow\infty$ and use DCT on $\lim_{T\uparrow\infty}\mathbb{E}[T^{r}Y_{T}\,|\,\mathscr{F}_{s}]$ and the right hand side of (A.3), we get
$\lim_{T\uparrow\infty}T^{r}\cdot Y_{T}=\liminf_{s\uparrow\infty}~\mathbb{E}\left[\lim_{T\uparrow\infty}T^{r}\cdot Y_{T}\bigg|\mathscr{F}_{s}\right]\leq\lim_{s\uparrow\infty}\sum_{t=s}^{\infty}\left((\alpha_{t}^{1+\gamma}\cdot t^{r}+t^{r-1})\cdot Y_{t}+t^{-p(1+\gamma)+r}Z_{t}\right)=0.$
This completes the proof. ∎
Appendix B Gronwall-type Inequalities
Proposition B.1.
Let $\{X_{n}\in\mathbb{R}:n\in\mathbb{N}_{0}\}$ and $\{Y_{n}\in\mathbb{R}:n\in\mathbb{N}_{0}\}$ be sequences, $n_{0}\in\mathbb{N}_{0}$, and $\beta>0$. The sequence
$Y_{n+n_{0}}=\beta^{2n}y+\mathbbm{1}_{\mathbb{N}}(n)\sum_{k=0}^{n-1}\beta^{2k}X_{n_{0}+n-1-k}~~,~~n\geq 0$
is the unique solution to the difference equation
$\begin{cases}Y_{n_{0}+n}=\beta^{2}\,Y_{n_{0}+n-1}+X_{n_{0}+n-1}~,~n>0.\\ Y_{n_{0}}=y.\end{cases}$ (B.1)
If $Z_{n}$ satisfies $Z_{n_{0}+n}\leq\beta^{2}\,Z_{n_{0}+n-1}+X_{n_{0}+n-1}$ (resp. $\geq$) and $Z_{n_{0}}\leq y$ (resp. $\geq$), then $Z_{n_{0}+n}\leq Y_{n_{0}+n}$ (resp. $\geq$).
Proof.
We have $Y_{n_{0}}=\beta^{0}y=y$ and $Y_{n_{0}+1}=\beta^{2}y+X_{n_{0}}=\beta^{2}Y_{n_{0}}+X_{n_{0}}$. Moreover, for $n>0$,
$X_{n_{0}+n}+\beta^{2}Y_{n_{0}+n}$ $=X_{n_{0}+n}+\beta^{2(n+1)}y+\sum_{k=0}^{n-1}\beta^{2k+2}X_{n_{0}+n-1-k}$
$=X_{n_{0}+n}+\beta^{2(n+1)}y+\sum_{k=1}^{n}\beta^{2k}X_{n_{0}+n-k}$
$=Y_{n_{0}+n+1}.$
For uniqueness, if $Z_{n_{0}+n}$ satisfies (B.1), then $Z_{n_{0}+n}-Y_{n_{0}+n}=\beta^{2}(Z_{n_{0}+n-1}-Y_{n_{0}+n-1})$ for all $n>0$ with $Z_{n_{0}}-Y_{n_{0}}=0$. This implies $Z_{n_{0}+n}-Y_{n_{0}+n}=(Z_{n_{0}}-Y_{n_{0}})\,\beta^{2n}=0$ for all $n\geq 0$.
For the second part, we show by induction that $Z_{n_{0}+n}\leq Y_{n_{0}+n}$. Note that $Z_{n_{0}}\leq y=Y_{n_{0}}$ and $Z_{n_{0}+1}\leq\beta^{2}y+X_{n_{0}}=Y_{n_{0}+1}$. For the inductive step, assume $Z_{n_{0}+n}\leq Y_{n_{0}+n}$ for $n\leq N$. From the first claim, $Y_{n_{0}+n+1}=\beta^{2}Y_{n_{0}+n}+X_{n_{0}+n}$. Therefore, we have
$Z_{n_{0}+N+1}\leq\beta^{2}\,Z_{n_{0}+N}+X_{n_{0}+N}\leq\beta^{2}\,Y_{n_{0}+N}+X_{n_{0}+N}=Y_{n_{0}+N+1}.$
On the other hand, if $Z_{n_{0}+n}\geq\beta^{2}Z_{n_{0}+n-1}+X_{n_{0}+n}$, then we may consider
$\begin{cases}-Z_{n_{0}+n}\leq\beta^{2}(-Z_{n_{0}+n-1})+(-X_{n_{0}+n-1})\\ -Z_{n_{0}}\leq-y.\end{cases}$
We can apply the previous claim and get $Z_{n_{0}+n}\leq-Y_{n_{0}+n}$. ∎
Proposition B.2.
Let $A\geq 0$, $X_{n}\geq 0$, $c_{n}\geq 0$.
$X_{n}\leq A+\sum_{k=1}^{n}c_{k-1}\,X_{k-1}~~~,~~~X_{0}\leq A~~~\implies X_{n}\leq A\prod_{k=1}^{n}\left(1+c_{k-1}\right).$
Proof.
Define the sequence
$Y_{n}=A+\sum_{k=1}^{n}c_{k-1}\,X_{k-1}~~~,~~~Y_{0}=A$
Observe that $Y_{n}-Y_{n-1}=c_{n-1}X_{n-1}\leq c_{n-1}Y_{n-1}$ for all $n\geq 1$, which implies
$\frac{Y_{n}}{Y_{n-1}}\leq 1+c_{n-1}~~,~~\forall n\geq 1.$
The above is a telescoping product. Along with the assumption $X_{n}\leq Y_{n}$, the above argument yields the desired result. ∎
Appendix C Some Facts Used in Proofs
Proposition C.1.
If the first two properties in assumption 2.1 are both satisfied, then $F$ is convex and $(\gamma,L)$-smooth.
Proof.
By the convexity of $\ell(z,\cdot)$, we get $\ell(z,a\,w_{1}+(1-a)w_{2})\leq a\,\ell(z,w_{1})+(1-a)\ell(z,w_{2})$. Applying the expectation w.r.t $\rho$ proves that $F$ is convex. Now, observe that we have
$\|\nabla F(w_{1})-\nabla F(w_{2})\|$ $=\|\mathbb{E}_{\rho}[\nabla\ell(Z,w_{1})-\nabla\ell(Z,w_{2})]\|$
$\leq\mathbb{E}_{\rho}\|\nabla\ell(Z,w_{1})-\nabla\ell(Z,w_{2})\|$
$\leq L\,\|w_{1}-w_{2}\|^{\gamma}.$
The proof is complete. ∎
Proposition C.2.
If $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $(\gamma,L)$-smooth, then (2.1) holds.
Proof.
Let $\phi(t)=f(x+t(y-x))$. Note that $\phi$ is continuously differentiable. Using the fundamental theorem of calculus on $\phi$ yields
$f(y)$ $=f(x)+\int_{0}^{1}\langle\nabla f(x+t(y-x))-\nabla f(x),\,y-x\rangle\,dt$
$\leq f(x)+\int_{0}^{1}\|\nabla f(x+t(y-x))-\nabla f(x)\|\,\|y-x\|\,dt$
$\leq f(x)+L\,\|y-x\|^{1+\gamma}\int_{0}^{1}t^{\gamma}\,dt.$
Integrating the right hand side yields the desired result. ∎
Lemma C.3.
([12, lemma 1]) Let $b_{t}$, $\eta_{t}$ be two non-negative sequences and $a_{t}$ be a sequence of vectors in a vector space $X$. Let $p\geq 1$ and assume $\sum_{t=1}^{\infty}\eta_{t}b_{t}^{p}<\infty$ and $\sum_{t=1}^{\infty}\eta_{t}=\infty$. Assume also that there exists $L>0$ such that both of the following conditions hold
  1. 1.
    $|b_{t+\tau}-b_{t}|\leq L\left(\sum_{i=t}^{t+\tau}\eta_{i}b_{i}+\left\|\sum_{i=t}^{t+\tau-1}\eta_{i}a_{i}\right\|\right)$.
  2. 2.
    $\left\|\sum_{i=1}^{\infty}\eta_{i}a_{i}\right\|<\infty$.
Then $b_{t}\rightarrow 0$.
Lemma C.4.
([8, lemma 13]). If $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $(\gamma,L)$-smooth and convex, then
$\frac{\gamma\,\|\nabla f(y)-\nabla f(x)\|^{\frac{1+\gamma}{\gamma}}}{(1+\gamma)L^{1/\gamma}}\leq f(y)-f(x)-\langle\nabla f(x),\,y-x\rangle\leq\frac{L\,\|y-x\|^{1+\gamma}}{1+\gamma}.$
Lemma C.5.
([8, lemma 14]). If assumption 2.1 holds and $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$, then for all $\theta\in(0,1]$,
$\mathbb{E}_{\rho}[\|\nabla\ell(Z,w)\|^{1+\theta}]\leq 2^{\theta}L^{1/\gamma}(1+\theta)\,[F(w)-F_{*}]+\frac{2^{\theta}(1-\gamma\theta)}{1+\gamma}+2^{\theta}\,\mathbb{E}_{\rho}[\|\nabla\ell(Z,w_{*})\|^{1+\theta}].$
Lemma C.6.
Assume assumption 2.1 holds and $\ell(z,\cdot)$ is convex for all $z\in\pazocal{Z}$.
  1. 1.
    There exists $a_{0},a_{1}\in\mathbb{R}_{+}$ such that $\|\nabla F(w)\|^{2}\leq a_{1}\,F(w)+a_{0}$ where $a_{1}$ is in (3a).
  2. 2.
    $F(w_{t})\leq F_{*}+\frac{L}{1+\gamma}\|w_{t}-w_{*}\|^{1+\gamma}$.
  3. 3.
    If in addition assumption 2.3 holds, then there exists $a_{1},a_{2},a_{4},a_{5},a_{6},a_{7}\in\mathbb{R}_{+}$ such that
    1. (a)
      $\|\nabla\ell(z,w)\|^{2}\leq a_{1}\,\ell(z,w)+a_{2}$.
    2. (b)
      $\|\nabla F(w)-\nabla\ell(z,w)\|^{2}\leq 6L^{2}\,\|w-w_{*}\|^{2\gamma}+a_{4}$.
    3. (c)
      $\|\nabla\ell(z,w)\|^{2}\leq 2L^{2}\,\|w-w_{*}\|^{2\gamma}+a_{5}$.
    4. (d)
      $\mathbb{E}[\|\nabla F(w_{t})-\nabla\ell(Z,w_{t})\|^{2}\,|\,\mathscr{F}_{t}]\leq a_{6}\,(F(w_{t})-F_{*})+a_{7}$.
Proof.
Proof of (1). We use the same argument as (C.1), but replace $\ell(z,w_{*})$ with $F_{*}$.
Proof of (2). Use $F(w_{t})=F(w_{t})-F_{*}+F_{*}$ and apply lemma C.4.
Proof of (3.a). Applying lemma C.4 and subsequently Young’s inequality [8, eq 4.3] with $p=(1+\gamma)/(1-\gamma)$ and $q=(1+\gamma)/(2\gamma)$ for $\gamma\in(0,1]$ yields
$\|\nabla\ell(z,w)\|^{2}$ $\leq\left(\frac{1+\gamma}{\gamma}\right)^{\frac{2\gamma}{1+\gamma}}L^{\frac{2}{1+\gamma}}\left(\ell(z,w)-\ell(z,w_{*})\right)^{\frac{2\gamma}{1+\gamma}}$ (C.1)
$\leq\frac{1-\gamma}{1+\gamma}\left(\frac{1+\gamma}{\gamma}\cdot L^{1/\gamma}\right)^{\frac{2\gamma}{1-\gamma}}+\frac{2\gamma}{1+\gamma}\left(\ell(z,w)-\ell(z,w_{*})\right)$
$\leq\frac{2\gamma}{1+\gamma}\,\ell(z,w)+\frac{1-\gamma}{1+\gamma}\left(\frac{1+\gamma}{\gamma}\cdot L^{1/\gamma}\right)^{\frac{2\gamma}{1-\gamma}}+\frac{2\gamma}{1+\gamma}\cdot\sup_{z\in\pazocal{Z}}\ell(z,w_{*}).$
Proof of (3.b). Observe that if assumption 2.3 and a minimizer $w_{*}\in\pazocal{W}$ exists and $\ell(z,\cdot)$ is convex in the second slot, then applying (3a) yields
$\|\nabla F(w)-\nabla\ell(z,w)\|^{2}$ $\leq 2\,\|\nabla F(w)\|^{2}+2\,\|\nabla\ell(z,w)\|^{2}$
$\leq 2L^{2}\,\|w_{t}-w_{*}\|^{2\gamma}+4\,\|\nabla\ell(z,w_{*})\|^{2}+4\|\nabla\ell(z,w)-\nabla\ell(z,w_{*})\|^{2}$
$\leq 6L^{2}\|w_{t}-w_{*}\|^{2\gamma}+4\,a_{1}\,\sup_{z\in\pazocal{Z}}\ell(z,w_{*})+4\,a_{2}.$
Proof of (3.c). By assumption 2.3 and claim (1), we get
$\|\nabla\ell(z,w)\|^{2}$ $\leq 2\|\nabla\ell(z,w_{*})\|^{2}+2\|\nabla\ell(z,w)-\nabla\ell(z,w_{*})\|^{2}$
$\leq\left(2a_{1}\sup_{z\in\pazocal{Z}}\ell(z,w_{*})+2a_{2}\right)+2L^{2}\|w-w_{*}\|^{2\gamma}$
Proof of (3.d). Another way of estimating $\|\nabla F(w)-\nabla\ell(z,w)\|^{2}$ is by using (1)
$\|\nabla F(w)-\nabla\ell(z,w)\|^{2}$ $\leq 2\|\nabla F(w_{t})\|^{2}+2\|\nabla\ell(Z_{t},w_{t})\|^{2}$
$\leq 2a_{1}\,(F(w)-F_{*})+2a_{1}F_{*}+2a_{0}+2a_{1}\,\ell(z,w)+2a_{2}.$
Applying $\mathbb{E}[\cdot\,|\,\mathscr{F}_{t}]$ yields
$\mathbb{E}[\|\delta m_{t}\|^{2}\,|\,\mathscr{F}_{t}]$ $\leq 2a_{1}\,(F(w_{t})-F_{*})+2a_{1}\,F(w_{t})+2(a_{0}+a_{2})$
$=4a_{1}\,(F(w_{t})-F_{*})+2(a_{0}+a_{2}+a_{1}\,F_{*}).$
This completes the proof. ∎
Proposition C.7.
Let $\delta m_{t}$ be as in (1.2) corresponding to the SHB iterates $w_{t}$ in (1.1).
  1. 1.
    $\mathbb{E}[\delta m_{t}\,|\,\mathscr{F}_{s}]=\mathbf{0}$ and $\mathbb{E}[\delta m_{s}^{(i)}\,\delta m_{t}^{(j)}]=0$ for all $s<t$ and $i\neq j$.
  2. 2.
    $\mathbb{E}[\langle\nabla F(w_{t}),\,\delta m_{t}\rangle\,|\,\mathscr{F}_{t}]=0$ for all $t>0$.
  3. 3.
    If $\alpha_{t}$ is a sequence in $\mathbb{R}$, then $\mathbb{E}\left\|\sum_{s=t}^{T}\alpha_{s}\,\delta m_{s}\right\|^{2}=\sum_{s=t}^{T}\alpha_{s}^{2}\,\mathbb{E}\|\delta m_{s}\|^{2}$ for all $t\leq T$.
Proof.
By (2.3), we have $\mathbb{E}[\delta m_{t}\,|\,\mathscr{F}_{t}]=\nabla F(w_{t})-\mathbb{E}[\nabla\ell(Z,w_{t})\,|\,\mathscr{F}_{t}]=0$. Furthermore,
$\mathbb{E}[\delta m_{t}\,|\,\mathscr{F}_{t-1}]$ $=\mathbb{E}[\nabla F(w_{t})\,|\,\mathscr{F}_{t-1}]-\mathbb{E}[\nabla\mathbb{E}_{\rho}[\ell(Z,w_{t})]\,|\,\mathscr{F}_{t-1}]=0.$
$\mathbb{E}[\delta m_{s}^{(i)}\,\delta m_{t}^{(j)}]$ $=\mathbb{E}[\delta m_{s}^{(i)}\,\mathbb{E}[\delta m_{t}^{(j)}\,|\,\mathscr{F}_{t-1}\,|\,\mathscr{F}_{s}]]=0~~,~~s<t.$
Since the standard inner product on $\mathbb{R}^{d}$ is a finite sum and the conditional expectation is linear, applying the first claim yields
$\mathbb{E}[\langle\nabla F(w_{t}),\,\delta m_{t}\rangle\,|\,\mathscr{F}_{t}]=\sum_{i=1}^{d}\mathbb{E}[\nabla F(w_{t})^{(i)}\delta m_{t}^{(i)}\,|\,\mathscr{F}_{t}]=\langle\nabla F(w_{t}),\,\mathbb{E}[\delta m_{t}\,|\,\mathscr{F}_{t}]\rangle=0.$
For the last claim, we have by the first claim,
$\mathbb{E}\left\|\sum_{s=t}^{T}\alpha_{s}\,\delta m_{t}\right\|^{2}$ $=\sum_{r,s=t}^{T}\alpha_{r}\cdot\alpha_{s}\cdot\mathbb{E}\langle\delta m_{r},\,\delta m_{s}\rangle$
$=\left(\sum_{r=s=t}^{T}\alpha_{s}^{2}\,\mathbb{E}\|\delta m_{s}\|^{2}\right)+2\sum_{r<s=t}^{T}\alpha_{r}\cdot\alpha_{s}\cdot\mathbb{E}\langle\delta m_{r},\,\delta m_{s}\rangle$
$=\sum_{r=s=t}^{T}\alpha_{s}^{2}\,\mathbb{E}\|\delta m_{s}\|^{2}.$
This completes the proof. ∎